### v1

# On the splitting of weak nearly $\cal C$}-manifolds

Abstract
Weak metric $f$-manifolds, i.e., the linear complex structure on the contact distribution is replaced by a nonsingular skew-symmetric tensor, defined by V. Rovenski and R. Wolak (2022), allow a new look at the theory of almost contact and metric $f$-manifolds. In this paper, we study the curvature and topology of a structure of this type, called the weak nearly $\cal C$}-structure; specifically, we are guided by the following question: Under what conditions are weak nearly $\cal C$-manifolds locally the Riemannian products? Our theorems generalize results of H. Endo (2005), A. Nicola–G. Dileo–I. Yudin (2018) on nearly cosymplectic manifolds and V. Rovenski (2024) on weak nearly cosymplectic manifolds to the context of weak nearly $\cal C$}-manifolds. Namely, we find conditions under which a $(2n+s)$-dimensional weak nearly $\cal C$}-manifold becomes a Riemannian product, and characterize $(4+s)$-dimensional weak nearly $\cal C$}-manifolds. The consequences of these theorems present new results for nearly $\cal C$}-manifolds.
Keywords: weak nearly $\cal C$}-manifold, weak nearly Kähler manifold, curvature tensor, foliation, Killing vector field.
Mathematics Subject Classifications (2010) 53C15, 53C25, 53D15
1 Introduction
A. Gray introduced in [10] a subclass of almost Hermitian manifolds $(M^{2n},J,\bar{g})$, called nearly Kähler manifolds, using the condition that the symmetric part of $\bar{\nabla}J$ vanishes. D. Blair and D. Showers defined in [5] a subclass of almost contact metric manifolds $(M^{2n+1},{f},\xi,\eta,g)$, called nearly cosymplectic manifolds, using a similar condition that only the symmetric part of $\nabla{f}$ vanishes. These odd-dimensional counterparts of nearly Kähler manifolds are useful in classifying almost contact metric manifolds and have applications in physics [11]. In dimensions greater than 5, a nearly cosymplectic manifold is locally the Riemannian product ${\mathbb{R}}\times\bar{M}^{2n}$ or $B^{5}\times\bar{M}^{2n-4}$, where $\bar{M}$ is a nearly Kähler manifold and $B$ is a nearly cosymplectic manifold, see [14]. Any 5-dimensional nearly cosymplectic manifold has an Einstein metric of positive scalar curvature. For example, the sphere $\mathbb{S}^{5}$ is endowed with a nearly cosymplectic structure induced by the almost Hermitian structure of $\mathbb{S}^{6}$.
The $f$-structure introduced by K. Yano, see [21], on a smooth manifold $M^{2n+s}$ serves as a higher-dimensional analog of almost complex structures ($s=0$) and almost contact structures ($s=1$). The recent interest of mathematicians in the $f$-structure is motivated by the study of the dynamics and integration on contact foliations, $\mathfrak{g}$-foliations, and $s$-cosymplectic manifolds, see [1, 7, 9, 13]. Contact foliations generalize to higher dimensions the flow of the Reeb vector field on contact manifolds, and $\cal K$}-structures are a particular case of uniform $s$-contact structures. An important class of metric $f$-manifolds $(M^{\,2n+s},{f},\vec{\xi},\vec{\eta},g)$, where $\vec{\xi}=(\xi_{1},\ldots,\xi_{s})$ are orthonormal vector fields, $\vec{\eta}=(\eta^{1},\ldots,\eta^{s})$ are dual 1-forms, is given by almost $\cal C$}-manifolds, i.e., $\nabla{f}=0$, see [3]. Any such manifold is locally the product of $\mathbb{R}^{s}$ and a Kähler manifold. Nearly $\cal C$}-manifolds are defined similarly to nearly Kähler manifolds, starting from almost $\cal C$}-manifolds, that is by condition $(\nabla_{X}f)X=0\ (X\in TM)$, e.g., [2]. The question arises: Under what conditions are nearly $\cal C$-manifolds locally the Riemannian products?
In [15, 16], Rovenski et al. introduced metric structures on a smooth manifold that generalize the almost Hermitian structure, almost contact metric structure, and metric $f$-structure. These so-called “weak” metric structures (the linear complex structure on the contact distribution is approximated by a nonsingular skew-symmetric tensor) made it possible to take a new look at the classical structures and find new applications in differential geometry and physics [22]. V. Rovenski defined the weak nearly cosymplectic and weak nearly Kähler structures (generalizing nearly cosymplectic and nearly Kähler structures, respectively), and under certain conditions obtained their decomposition, see [18], generalizing the classical result in [14]. In [20], V. Rovenski initiated the study of a special class of metric $f$-manifolds called weak nearly $\cal C$-manifolds, which generalize nearly $\cal C$-manifolds. Guided by [18, 20], in the paper we investigate the above question for the case of weak nearly $\cal C$-manifolds.
In this article, we study the curvature and topology of weak nearly $\cal C$}-manifolds and find conditions (9) and (13) that are trivially satisfied by almost metric $f$-manifolds and under which weak nearly $\cal C$}-manifolds are locally Riemannian products. In Section 2, following the introductory Section 1, we recall necessary results on weak metric $f$-structures. In Section 3, we formulate auxiliary lemmas on the extrinsic geometry and curvature of weak nearly $\cal C$}- and weak nearly Kähler manifolds.
[…]
2 Basic concepts and examples of $f$-structure
In this section, we will review some well-known concepts and results, see [16, 17, 19, 20].
Recall that nearly Kähler manifolds $(M,J,g)$ are defined by A. Gray [10] using the condition that only the symmetric part of $\nabla J$ vanishes, in contrast to the Kähler case where $\nabla J=0$.
Definition 1.
A Riemannian manifold $(M,g)$ of even dimension equipped with a skew-symmetric (1,1)-tensor f{f} such that the tensor ${f}^{\,2}$ is negative-definite is called a weak Hermitian manifold. Such (M,f,g)(M,{f},g) is called a weak Kähler manifold if $\nabla{f}=0$, where $\nabla$ is the Levi-Civita connection. A weak Hermitian manifold is called a weak nearly Kähler manifold, if $(\nabla_{X}{f})Y+(\nabla_{Y}{f})X=0\ (X,Y\in\mathfrak{X}_{M})$.
Definition 2.
weak metric $f$-structure on a smooth manifold $M^{2n+s}$ $(n,s>0)$ is a set $({f},Q,\vec{\xi},\vec{\eta}$, $g)$, where f{f} is a skew-symmetric $(1,1)$-tensor of rank $2n$, $Q$ is a self-adjoint nonsingular $(1,1)$-tensor, $\vec{\xi}=(\xi_{1},\ldots,\xi_{s})$ are orthonormal vector fields, $\vec{\eta}=(\eta^{1},\ldots,\eta^{s})$ are dual 1-forms, and $g$ is a Riemannian metric on $M$, satisfying
${f}^{2}=-Q+\sum\nolimits_{i=1}^{s}{\eta^{i}}\otimes{\xi_{i}},\quad{\eta^{i}}({\xi_{j}})=\delta^{i}_{j},\quad Q\,{\xi_{i}}={\xi_{i}},$ (1)
$g({f}X,{f}Y)=g(X,Q\,Y)-\sum\nolimits_{i=1}^{s}{\eta^{i}}(X)\,{\eta^{i}}(Y)\quad(X,Y\in\mathfrak{X}_{M}).$ (2)
In this case, $(M^{2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ is called a weak metric $f$-manifold.
Putting $Y=\xi_{j}$ in (2), and using ${\eta^{i}}({\xi_{j}})=\delta^{i}_{j}$, we get
$\eta^{j}(X)=g(X,\xi_{j});$ (3)
thus, $\xi_{j}$} is orthogonal to the contact distribution ${\cal D}:=\bigcap_{\,i=1}^{s}\ker{\eta^{i}}$. For a more intuitive understanding of the role of $Q$ in the metric $f$-structure, we explain the following properties:
${f}\,{\xi_{i}}=0,\quad{\eta^{i}}\circ{f}=0,\quad\eta^{i}\circ Q=\eta^{i},\quad[Q,\,{f}]=0.$
By (1), $f^{2}\xi_{i}=0$ is true. From this and (1), we get $f^{3}+fQ=0$. By this, $Q\xi_{i}=\xi_{i}$ and $f^{2}\xi_{i}=0$ we get $0=-f^{3}\xi_{i}=fQ\xi_{i}=f\xi_{i}$. By $f\xi_{i}=0$, (3), and the skew-symmetry of $f$, we get $\eta^{i}(fX)=g(fX,\xi_{i})=-g(X,f\xi_{i})=0$. From this and condition ${\rm rank}\,f=2n$, we conclude that $f$ the distribution $\cal D$} of a weak metric $f$-structure is f{f}-invariant, ${\cal D}=f(TM)$ and $\dim{\cal D}=2n$. By this and $f^{3}+fQ=0$, we get $f^{3}X=f^{2}(fX)=-QfX$; hence, $f^{3}+Qf=0$. This and $f^{3}+fQ=0$ yield $fQ=Qf$. By symmetry of $Q$ and $Q\xi_{i}=\xi_{i}$, we get $\eta^{i}(QX)=g(QX,\xi_{i})=g(X,Q\xi_{i})=g(X,\xi_{i})=\eta^{i}(X)$. Therefore, the tangent bundle splits, $TM={\cal D}\oplus\ker f$ as the complementary orthogonal sum of its subbundles, the contact distribution $\cal D$}, and $\ker f=\operatorname{span}\{\xi_{1},\dots,\xi_{s}\}$ called the Reeb distribution.
The exterior derivative of a 1-form $\omega$} is given by
$d\omega(X,Y)=X({\omega}(Y))-Y({\omega}(X))-{\omega}([X,Y])\quad(X,Y\in\mathfrak{X}_{M}).$
The fundamental $2$-form $\Phi$ on $(M^{2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ is given by $\Phi(X,Y)=g(X,{f}Y)$ for all $X,Y\in\mathfrak{X}_{M}$. Recall the co-boundary formula for exterior derivative $d$ on a $2$-form $\Phi$,
$d\Phi(X,Y,Z)$ $=X\,\Phi(Y,Z)+Y\,\Phi(Z,X)+Z\,\Phi(X,Y)$
$\ -\Phi([X,Y],Z)-\Phi([Z,X],Y)-\Phi([Y,Z],X).$
Note that the above equality yields
$d\Phi(X,Y,Z)=(\nabla_{X}\,\Phi)(Y,Z)+(\nabla_{Y}\,\Phi)(Z,X)+(\nabla_{Z}\,\Phi)(X,Y).$ (4)
A weak metric $f$-structure is called a weak $f$-K-contact structure if $\Phi=d{\eta^{1}}=\ldots=d{\eta^{s}}$ is valid and all characteristic vector fields $\xi_{i}$} are Killing, i.e. $(\pounds_{{\xi_{i}}}\,g)(X,Y)=g(\nabla_{X}{\xi_{i}},Y)+g(\nabla_{Y}{\xi_{i}},X)=0$.
Another important case of a weak metric $f$-structure is a weak almost $\cal K$}-structure, that is $d\Phi=0$. Its special case, is a weak almost $\cal C$}-structure, that is $\Phi$ and $\eta^{i}\ (1\leq i\leq s)$ are closed forms. For $s=1$, weak almost $\cal C$}- manifolds reduce to weak almost cosymplectic manifolds. For weak almost $\cal K$}-manifolds (and their subclass, weak almost $\cal C$}-manifolds), the Reeb distribution $\ker f$ is tangent to a foliation. Moreover, weak almost $\cal C$}-manifolds satisfy the following conditions:
$[\xi_{i},\xi_{j}]$ $=0\quad(1\leq i,j\leq s),$ (5)
$g(\nabla_{X}\,\xi_{i},\ \xi_{j})$ $=0\quad(X\in\mathfrak{X}_{M},\ 1\leq i,j\leq s),$ (6)
trivial for $s=1$. The following condition is a corollary of (6):
$\eta^{k}(\nabla_{\xi_{i}}\,\xi_{j})=0\quad(1\leq i,j,k\leq s).$ (7)
A metric $f$-structure, satisfying (5) and $\nabla f=0$, is a weak almost $\cal C$}-structure. We consider the case when only the symmetric part of $\nabla f$ vanishes.
Definition 3.
A weak metric $f$-structure is called a weak nearly $\cal C$}-structure if
$(\nabla_{X}{f})Y+(\nabla_{Y}{f})X=0.$ (8)
Taking derivative of g(fV,Z)=g(V,fZ)g({f}V,Z)=-g(V,{f}Z), we see that $\nabla_{Y}{f}$ of a weak nearly $\cal C$}-manifold is skew-symmetric: $g((\nabla_{Y}{f})V,Z)=-g((\nabla_{Y}{f})Z,V)$. Taking the derivative of this, we see that $\nabla^{2}_{X,Y}{f}$ is also skew-symmetric: $g((\nabla^{2}_{X,Y}{f})V,Z)=-g((\nabla^{2}_{X,Y}{f})Z,V)$, where $\nabla^{2}_{X,Y}:=\nabla_{X}\nabla_{Y}-\nabla_{\nabla_{X}Y}$.
Example 1.
$(i)$ Any weak Kähler manifold is obviously weak nearly Kähler. Take two or more (nearly) Kähler manifolds $(M_{j},g_{j},J_{j})$, where $J_{j}^{2}=-{\rm Id}_{\,j}$. The product $\prod_{j}(M_{j},g_{j},\sqrt{\lambda_{j}}J_{j})$, where $\lambda_{j}\neq 1$ are different positive constants, is a weak (nearly) Kähler manifold with $Q=\bigoplus_{j}\lambda_{j}{\rm Id}_{j}$. The classification of weak nearly Kähler manifolds in dimensions $\geq 4$ is an open problem.
$(ii)$ To construct a weak metric $f$-structure $({f},Q,\vec{\xi},\vec{\eta},g)$ on the Riemannian product $M=\bar{M}\times\mathbb{R}^{s}$ of a weak almost Hermitian manifold $(\bar{M},\bar{f},\bar{g})$ with $\Omega(X,Y)=\bar{g}(X,\bar{f}Y)$ and a Euclidean space $(\mathbb{R}^{s},dy^{2})$, we take any point $(x,y)$ of $M$ and set
$\xi_{i}=(0,\partial_{\,y^{i}}),\ \ \eta^{i}=(0,dy^{i}),\ \ {f}(X,\partial_{\,y^{i}})=(\bar{f}X,0),\ \ Q(X,\partial_{\,y^{i}})=(-\bar{f}^{\,2}X,\partial_{\,y^{i}}),$
where $X\in T_{x}\bar{M}$. Note that $\nabla f=0$ if and only if $\overline{\nabla}\bar{f}=0$. On the other hand, $\overline{\nabla}\bar{f}=0$ if and only if $d\Omega=0$, see (4) with $\Phi=\Omega$, i.e., $(\bar{M},\Omega)$ is a symplectic manifold.
$(iii)$ To construct a weak nearly $\cal C$-structure $({f},Q,\vec{\xi},\vec{\eta},g)$ on the Riemannian product $M=\bar{M}\times\mathbb{R}^{s}$ of a weak nearly Kähler manifold $(\bar{M},\bar{f},\bar{g})$ and a Euclidean space $(\mathbb{R}^{s},dy^{2})$, we take any point $(x,y)$ of $M$ and set, as in Example 1(ii)(ii),
$\xi_{i}=(0,\partial_{\,y^{i}}),\ \ \eta^{i}=(0,dy^{i}),\ \ {f}(X,\partial_{\,y^{i}})=(\bar{f}X,0),\ \ Q(X,\partial_{\,y^{i}})=(-\bar{f}^{\,2}X,\partial_{\,y^{i}}).$
By (5), the Reeb distribution $\ker f$ of weak almost $\cal C$}-manifolds is tangent to a $\mathfrak{g}$-foliation with an abelian Lie algebra. A foliation of dimension $s$ on a smooth connected manifold $M$ is called a $\mathfrak{g}$-foliation [1] (where $\mathfrak{g}$ is a Lie algebra of dimension $s$) if there exist complete vector fields $\xi_{1},\ldots,\xi_{s}$ on $M$ which, when restricted to each leaf, form a parallelism of this submanifold with a Lie algebra isomorphic to $\mathfrak{g}$. Below, we present an interesting way to construct a weak nearly $\cal C$}-structure on a manifold $M$ that has two nearly $\cal C$}-structures with the same Reeb vector fields.
Example 2.
Let a Riemannian manifold $(M^{2n+s},g)$ admit two nearly $\mathcal{C}$}-structures with common Reeb vector fields $\xi_{i}$ and one-forms $\eta^{i}=g(\xi_{i},\cdot)$ for $1\leq i\leq s$. Suppose that $f_{1}\neq f_{2}$ are such that the self-adjoint tensor $\psi:=f_{1}f_{2}+f_{2}f_{1}$ is not identically zero. Then the tensor $f:=(\cos t)\,f_{1}+(\sin t)\,f_{2}$ for small $t>0$ ($t$ is a constant) satisfies the nearly $\mathcal{C}$-condition (8). Indeed,
$(\nabla_{X}f)X$ $=\nabla_{X}(fX)-f(\nabla_{X}X)=\cos t\,(\nabla_{X}f_{1})X+\sin t\,(\nabla_{X}f_{2})X=0.$
Also, observe that $f^{2}=-\mathrm{id}+(\sin t\cos t)\,\psi+\sum\nolimits_{i=1}^{s}\eta^{i}\otimes\xi_{i}$. So, $Q=\mathrm{id}-(\sin t\cos t)\,\psi$, which is positive definite for small $t>0$. Thus, $(f,Q,\vec{\xi},\vec{\eta},g)$ is a weak nearly $\cal C$-structure on $M^{2n+s}$.
The following condition is trivial when $Q=\mathrm{id}$:
$(\nabla_{X}\,Q)Y=0\quad(X\in\mathfrak{X}_{M},\ Y\in{\cal D}).$ (9)
Using (9), we have
$(\nabla_{X}\,Q)Y=\sum\nolimits_{i=1}^{s}{\eta^{i}}(Y)(\nabla_{X}\,Q)\xi_{i}=-\sum\nolimits_{i=1}^{s}\eta^{i}(Y)\,\widetilde{Q}\nabla_{X}\,\xi_{i}\quad(X,Y\in\mathfrak{X}_{M}).$ (10)
Lemma 1.
On a weak nearly $\cal C$}-manifold satisfying (5) and (7), the Reeb distribution $\ker f$ defines a flat totally geodesic foliation; moreover, if (6) and (9) hold, then the vector fields $\xi_{i}$ are Killing.
For a Riemannian manifold $(M,g)$ equipped with a Killing vector field $\zeta$}, we get, see [21],
$\nabla_{X}\nabla_{Y}\,{\zeta}-\nabla_{\nabla_{X}Y}\,{\zeta}=R_{\,X,\,{\zeta}}\,Y,$ (11)
where $R_{{X},{Y}}=[\nabla_{X},\nabla_{Y}]-\nabla_{[X,Y]}$ is the curvature tensor. We have $g(R_{\,\xi,Z}\,{f}X,{f}Y)=0$ for a nearly cosymplectic manifold $(M^{2n+1},f,\xi,\eta,g)$, see [8]; thus its distribution ${\cal D}=\ker\eta$ is curvature invariant:
$R_{X,Y}Z\in{\cal D}\quad(X,Y,Z\in{\cal D}).$ (12)
For example, the property (12) is true for any distribution $\cal D$} on a real space form. A nearly $\cal C$}-manifold satisfying conditions (5) and (6), also satisfies (12) with ${\cal D}=\bigcap_{\,i=1}^{s}\ker{\eta^{i}}$, see Corollary 1. In our study of weak nearly $\cal C$}-manifolds, we assume a condition weaker than (12):
$R_{\widetilde{Q}X,Y}Z\in{\cal D}\quad(X,Y,Z\in{\cal D}).$ (13)
Obviously, conditions (9) and (13) become trivial for metric $f$-manifolds, i.e., when $Q=\mathrm{id}$.
3 Auxiliary results
In this section, we investigate weak nearly $\mathcal{C}$-manifolds satisfying the conditions (5), (6), (9) and (13), or only some of them. Namely, we prove several lemmas that do not require conditions (9) and (13) in the case of nearly $\mathcal{C}$-manifolds, i.e., $Q={\rm Id}$, and generalize to $s>1$ some well-known properties of weak nearly cosymplectic manifolds ($s=1$).
We define the tensor field ${\bf h}=(h_{1},\ldots,h_{s})$, as in the classical case, e.g., [8], where
$h_{i}=\nabla\xi_{i}\quad(1\leq i\leq s).$ (14)
As a consequence of Lemma 1, the tensor fields $h_{i}$ are skew-symmetric:
$g(h_{i}X,\,X)=g(\nabla_{X}\,\xi_{i},X)=\frac{1}{2}\,(\pounds_{\xi_{i}}\,g)(X,X)=0\quad(1\leq i\leq s).$
Using (6), we get $\eta^{j}\circ h_{i}=0$ and $h_{i}\,\xi_{j}=0$ for $1\leq i,j\leq s$. This shows that $\operatorname{dim(ker}h_{i})\geq 3$ for all $i$.
The following result generalizes [18, Lemma 3.1] (and Lemma 3.1 in [8]) for the case of $s>1$.
Lemma 2.
For a weak nearly $\cal C$}-manifold $(M^{\,2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ satisfying (5) and (6), we get
$(\nabla_{X}\,h_{i})\,\xi_{j}=-h_{i}h_{j}X\quad(1\leq i,j\leq s),$ (15)
$(\nabla_{X}{f})\,\xi_{i}=-{f}\,h_{i}X\quad(1\leq i\leq s),$ (16)
$h_{i}\,{f}+{f}\,h_{i}=0\quad(1\leq i\leq s)\quad(h_{i}\ {\rm anticommutes\ with}\ {f}),$ (17)
$h_{i}\,Q=Q\,h_{i}\quad(1\leq i\leq s)\quad(h_{i}\ {\rm commutes\ with}\ Q).$ (18)
Proof.
Differentiating the equality $h_{i}\,\xi_{j}=0$ and using (14), we obtain (15):
$0=\nabla_{X}\,(h_{i}\,\xi_{j})=(\nabla_{X}\,h_{i})\,\xi_{j}+h_{i}(\nabla_{X}\,\xi_{j})=(\nabla_{X}\,h_{i})\,\xi_{j}+h_{i}h_{j}X.$
Equation (16) is obtained by differentiating the relation $f\xi_{i}=0$ and using the definition of $h_{i}$. Differentiating $g({f}\,Y,\xi_{i})=0$ yields $0=g((\nabla_{X}{f})Y,\xi_{i})+g({f}\,Y,h_{i}X)$. Interchanging $X$ and $Y$ in the last equation, then adding the resulting expression to the original and applying (8), yields (17). Finally, from (1) and (6) we get $f^{2}h_{i}X-h_{i}f^{2}X=-Qh_{i}X+h_{i}QX$ for all $X\in\mathfrak{X}_{M}$ and (18) follows from this by applying (17). ∎
The following lemmas generalize certain formulas in [18, Lemmas 3.2-3.5] (and Lemmas 3.2 – 3.5 in [8]) for the case of $s>1$.
Lemma 3.
For a weak nearly $\cal C$}-manifold satisfying (9) we obtain
$g((\nabla_{X}{f}){f}Y,Z)$ $=g((\nabla_{X}{f})Y,{f}Z)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Y)g(h_{i}X,Z)+\eta^{i}(Z)g(h_{i}X,QY)\big],$ (19)
$g((\nabla_{{f}X}{f})Y,Z)$ $=g((\nabla_{X}{f})Y,{f}Z)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(X)g(h_{i}Z,Y)+\eta^{i}(Z)g(h_{i}X,QY)\big],$ (20)
$g((\nabla_{{f}X}{f}){f}Y,Z)$ $=g((\nabla_{X}{f})QZ,Y)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(X)g(h_{i}Z,{f}Y)+\eta^{i}(Y)g(h_{i}X,{f}Z)$
$\ -\eta^{i}(Z)g({f}h_{i}X,\widetilde{Q}Y)\big].$ (21)
Proof.
As in the proof of [8, Lemma 3.4], differentiating (1) and using (10), and the skew-symmetry of $\nabla_{X}{f}$, we get (19). We obtain (20) from (19) by the definition of weak nearly $\cal C$}-manifolds, followed by switching $X$ and $Y$. Replacing $Y$ by fY{f}Y in (20) and using (19) and (1), we get (3). ∎
Lemma 4.
The curvature tensor of a weak nearly $\cal C$}-manifold satisfies the equality
$g(R_{{f}X,Y}Z,V)+g(R_{X,{f}Y}Z,V)+g(R_{X,Y}{f}Z,V)+g(R_{X,Y}Z,{f}V)=0;$ (22)
moreover, if the conditions (9) and (13) are true, then
$R_{\,\xi_{j},\,\xi_{k}}=0\quad(1\leq j,k\leq s),$ (23)
$g(R_{\,\xi_{j},Z}\,{f}X,{f}Y)=0\quad(1\leq j\leq s)\quad(\mathcal{D}\ \mbox{\rm is curvature invariant}).$ (24)
Proof.
Differentiating (8), we find
$(\nabla^{2}_{X,Y}\,{f})Z+(\nabla^{2}_{X,Z}\,{f})Y=0\quad(X,Y,Z\in\mathfrak{X}_{M}).$ (25)
Applying the Ricci identity, see [8],
$g((\nabla^{2}_{X,Y}{f})V,Z)-g((\nabla^{2}_{Y,X}{f})V,Z)=g(R_{X,Y}{f}V,Z)+g(R_{X,Y}V,{f}Z),$ (26)
from (25) and the skew-symmetry of $\nabla^{2}_{X,Y}\,{f}$ we get
$g(R_{X,Y}Z,{f}V)-g(R_{X,Y}V,{f}Z)+g((\nabla^{2}_{X,Z}\,{f})Y,V)-g((\nabla^{2}_{Y,Z}\,{f})X,V)=0.$ (27)
By Bianchi and Ricci identities, we find
$g(R_{X,Y}Z,{f}V)$ $=-g(R_{Y,Z}X,{f}V)-g(R_{Z,X}Y,{f}V)$
$=g((\nabla^{2}_{Y,Z}\,{f})V,X)-g((\nabla^{2}_{Z,Y}\,{f})V,X)-g(R_{Y,Z}V,{f}X)-g(R_{Z,X}Y,{f}V).$ (28)
Substituting (3) into (27), we get that
$g(R_{X,Z}Y,{f}V)-g(R_{X,Y}V,{f}Z)-g(R_{Y,Z}V,{f}X)$
$-g((\nabla^{2}_{Z,Y}\,{f})V,X)-g((\nabla^{2}_{X,Z}\,{f})V,Y)=2\,g((\nabla^{2}_{Y,Z}\,{f})X,V).$ (29)
On the other hand, using the Ricci identity (26) again, we see that
$g(R_{X,Z}Y,{f}V)-g(R_{X,Z}V,{f}Y)-g((\nabla^{2}_{X,Z}\,{f})Y,V)+g((\nabla^{2}_{Z,X}\,{f})Y,V)=0.$ (30)
Adding (30) to (3) and applying (25), we get
$2\,g(R_{X,Z}Y,{f}V)-g(R_{X,Y}V,{f}Z)-g(R_{Y,Z}V,{f}X)-g(R_{X,Z}V,{f}Y)=2\,g((\nabla^{2}_{Y,V}\,{f})Z,X).$ (31)
Swapping $Y$ and $V$ in (31), we find
$2\,g(R_{X,Z}V,{f}Y)-g(R_{X,V}Y,{f}Z)-g(R_{V,Z}Y,{f}X)-g(R_{X,Z}Y,{f}V)=2\,g((\nabla^{2}_{V,Y}\,{f})Z,X).$ (32)
Subtracting (32) from (31), and using Bianchi and Ricci identities, we get the equality, which by replacing $Z$ and $Y$ gives (22).
Now, assume that (9) and (13) are true. Replacing $X$ by fX{f}X in (22) and using (1), we have
$-g(R_{QX,Y}Z,V)+\sum\nolimits_{i=1}^{s}\eta^{i}(X)\,g(R_{\,\xi_{i},Y}Z,V)+g(R_{{f}X,{f}Y}Z,V)$
$\quad+g(R_{{f}X,Y}{f}Z,V)+g(R_{{f}X,Y}Z,{f}V)=0.$ (33)
Exchanging $X$ and $Y$ in (3), we find
$g(R_{X,QY}Z,V)$ $+\sum\nolimits_{i=1}^{s}\eta^{i}(Y)\,g(R_{\,\xi_{i},X}Z,V)-g(R_{{f}X,{f}Y}Z,V)$
$+g(R_{{f}Y,X}{f}Z,V)+g(R_{{f}Y,X}Z,{f}V)=0.$ (34)
Subtracting (3) from (3), we obtain
$2\,g(R_{{f}X,{f}Y}Z,V)-2\,g(R_{X,Y}Z,V)+\sum\nolimits_{i=1}^{s}[\eta^{i}(X)\,g(R_{\,\xi_{i},Y}Z,V)-\eta^{i}(Y)\,g(R_{\,\xi_{i},X}Z,V)]$
$+g(R_{{f}X,Y}{f}Z,V)-g(R_{{f}Y,X}{f}Z,V)+g(R_{{f}X,Y}Z,{f}V)-g(R_{{f}Y,X}Z,{f}V)$
$-g(R_{\widetilde{Q}X,Y}Z,V)-g(R_{X,\widetilde{Q}Y}Z,V)=0.$ (35)
Then, replacing $Z$ by fZ{f}Z and also $V$ by fV{f}V in (22) and using (1), we get two equations
$g(R_{X,Y}QZ,V)$ $=\sum\nolimits_{i=1}^{s}\eta^{i}(Z)\,g(R_{X,Y}\,\xi_{i},V)+g(R_{X,Y}{f}Z,{f}V)$
$\quad+g(R_{X,{f}Y}{f}Z,V)+g(R_{{f}X,Y}{f}Z,V),$ (36)
$g(R_{X,Y}Z,QV)$ $=\sum\nolimits_{i=1}^{s}\eta^{i}(V)\,g(R_{X,Y}Z,\,\xi_{i})+g(R_{X,Y}{f}Z,{f}V)$
$\quad+g(R_{{f}X,Y}Z,{f}V)+g(R_{X,{f}Y}Z,{f}V).$ (37)
Adding (3) to (3), and substituting the gotten equation into (3), we have
$2\,g(R_{{f}X,{f}Y}Z,V)-2\,g(R_{X,Y}{f}Z,{f}V)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(X)\,g(R_{\,\xi_{i},Y}Z,V)-\eta^{i}(Z)\,g(R_{X,Y}\,\xi_{i},V)$
$\quad-\eta^{i}(V)\,g(R_{X,Y}Z,\,\xi_{i})\,-\eta^{i}(Y)\,g(R_{\,\xi_{i},X}Z,V)\big]+\delta(X,Y,Z,V)=0,$ (38)
where $\delta(X,Y,Z,V)=g(R_{X,Y}\widetilde{Q}Z,V)+g(R_{X,Y}Z,\widetilde{Q}V)-g(R_{\widetilde{Q}X,Y}Z,V)-g(R_{X,\widetilde{Q}Y}Z,V)$. Replacing $X$ by fX{f}X and also $Y$ by fY{f}Y in (3) and using (1), we obtain
$2\,g(R_{QX,QY}Z,V)+2\,\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Y)\,g(R_{\,\xi_{i},QX}Z,V)-\eta^{i}(X)\,g(R_{\,\xi_{i},QY}Z,V)\big]-2\,g(R_{{f}X,{f}Y}\,{f}Z,{f}V)$
$+\delta({f}X,{f}Y,Z,V)=\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Z)\,g(R_{\,\xi_{i},V}{f}X,{f}Y)-\eta^{i}(V)\,g(R_{\,\xi_{i},Z}\,{f}X,{f}Y)\big].$ (39)
Using $X=\xi_{j}$ and $Y=\xi_{k}$ in (3), we get $g(R_{\xi_{j},\xi_{k}}Z,V)=0$ for all $1\leq j,k\leq s$; hence (23). This also yields $g(R_{\xi_{j},\,\xi_{k}}fZ,fV)=0$. Taking $V=\xi_{j}$ in (3) and using the previous relation, we obtain
$2\,g(R_{\,QX,\,QY}Z,\xi_{j})+2\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Y)\,g(R_{\,\xi_{i},QX}Z,\xi_{j})-\eta^{i}(X)\,g(R_{\,\xi_{i},QY}Z,\xi_{j})\big]$
$+\,g(R_{\,\xi_{j},Z}\,{f}X,{f}Y)+\delta({f}X,{f}Y,Z,\,\xi_{j})=0\quad(1\leq j\leq s).$ (40)
Replacing $X$ by fX{f}X and $Y$ by fY{f}Y in (3), we obtain
$4\,g(R_{\,QfX,\,QfY}Z,\xi_{j})-2\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Y)\,g(R_{\,\xi_{i},QX}Z,\xi_{j})-\eta^{i}(X)\,g(R_{\,\xi_{i},QY}Z,\xi_{j})\big]$
$+2\,g(R_{\,\xi_{j},Z}\,{Q}X,{Q}Y)+2\,\delta({f}^{2}X,{f}^{2}Y,Z,\,\xi_{j})=0\quad(1\leq j\leq s).$
Adding this to (3) and using $Q=Id+\widetilde{Q}$, we get
$3\,g(R_{\,\xi_{j},Z}\,{f}X,{f}Y)=-4\,g(R_{\,\xi_{j},Z}\,\widetilde{Q}{f}X,\,\widetilde{Q}{f}Y)-4\,g(R_{\,\xi_{j},Z}\,{f}X,\,\widetilde{Q}{f}Y)$
$-4\,g(R_{\,\xi_{j},Z}\,\widetilde{Q}{f}X,\,{f}Y)+\delta({f}X,{f}Y,Z,\,\xi_{j})+2\,\delta({f}^{2}X,{f}^{2}Y,Z,\,\xi_{j})\quad(1\leq j\leq s).$ (41)
From (3), using (13), (23) and the equality $\delta({f}X,{f}Y,Z,\xi_{j})=0$, we get (24). Since ${f}|_{\mathcal{D}}$ is non-degenerate, the distribution $\mathcal{D}$ is curvature invariant, see (12). ∎
Corollary 1.
For a nearly $\cal C$}-manifold the contact distribution is curvature invariant, i.e., $R_{X,Y}Z\in\mathcal{D}$ for all $X,Y,Z\in\mathcal{D}$.
Proof.
For a nearly $\cal C$}-manifold, $Q={\rm Id}$ is valid, which guarantees that (9) and (13) are satisfied. Hence, by Lemma 4 and (24), we have $g(R_{\xi_{i},fZ}fX,fY)=0$ for all $1\leq i\leq s$ and $X,Y,Z\in\mathfrak{X}_{M}$. ∎
Lemma 5.
For a weak nearly $\cal C$}-manifold satisfying conditions (5), (6), (9) and (13), we get
$R_{\,\xi_{i},\,X}Y=-(\nabla_{X}h_{i})Y,$ (42)
$(\nabla_{X}h_{i})Y=\sum\nolimits_{j=1}^{s}[g(h_{i}h_{j}X,Y)\,\xi_{j}-\eta^{j}(Y)\,h_{i}h_{j}X],$ (43)
${\rm Ric}\,(\xi_{i},Z)=-\sum\nolimits_{j=1}^{s}\eta^{j}(Z)\,\operatorname{tr}(h_{i}h_{j}).$ (44)
In particular, $\nabla_{\xi_{j}}h_{i}=0$ and ${\rm tr}(h_{i}^{2})=const$ for all $1\leq i,j\leq s$. Moreover,
$g(R_{\,\xi_{i},X}Y,Z)$ $=-g((\nabla_{X}\,h_{i})Y,Z)=\sum\nolimits_{j=1}^{s}[\eta^{j}(Y)\,g(h_{i}h_{j}X,Z)-\eta^{j}(Z)\,g(h_{i}h_{j}X,Y)],$ (45)
$g(R_{{f}X,{f}Y}Z,V)$ $=g(R_{X,Y}{f}Z,{f}V)-\frac{1}{2}\,\delta(X,Y,Z,V),$ (46)
$g(R_{{f}X,{f}Y}{f}Z,{f}V)$ $=g(R_{QX,QY}Z,V)+\frac{1}{2}\,\delta({f}X,{f}Y,Z,V)$
$\quad+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Y)\,g(R_{\,\xi_{i},QX}Z,V)-\eta^{i}(X)\,g(R_{\,\xi_{i},QY}Z,V)\big].$ (47)
Proof.
By Lemma 1, $\xi_{i}$ is a Killing vector field, and applying (11) in conjunction with (14), we get (42). Replacing $Y$ by fY{f}Y and $Z$ by fZ{f}Z in $g(R_{\,\xi_{i},X}Y,Z)=-g((\nabla_{X}\,h_{i})Y,Z)$, see (42), and using (24), we get $g((\nabla_{X}h_{i})\,{f}Y,{f}Z)=0$, hence,
[…]
Then, using (48), we find the $\ker f$-component and $\cal D$}-component of $(\nabla_{X}h_{i})Y$:
$g((\nabla_{X}h_{i})Y,\,\xi_{j})=g(\nabla_{X}(h_{i}\,Y),\,\xi_{j})=-g(h_{i}\,Y,\,\nabla_{X}\,\xi_{j})=g(h_{i}h_{j}X,Y),$
$g((\nabla_{X}\,h_{i})Y,Z)=\sum\nolimits_{j=1}^{s}\eta^{j}(Y)\,g((\nabla_{X}h_{i})\,\xi_{j},\,Z)=-\sum\nolimits_{j=1}^{s}\eta^{j}(Y)\,g(h_{i}h_{j}X,Z)\quad(Z\in{\cal D}),$
from which (43) follows. From (43) with $X=\xi_{k}$ we find $\nabla_{\xi_{k}}h_{i}=0$. Note that for $Y=\xi_{j}$, (43) reduces to (15). By (43) and (42), we get (45). Using (3) and (45), we obtain
$2g(R_{{f}X,{f}Y}Z,V)-2g(R_{X,Y}{f}Z,{f}V)+\!\sum\nolimits_{i,j=1}^{s}\!\big[\eta^{i}(X)\eta^{j}(Z)g(h_{i}h_{j}Y,V)-\eta^{i}(Z)\eta^{j}(X)g(h_{i}h_{j}V,Y)$
$-\eta^{i}(X)\eta^{j}(V)g(h_{i}h_{j}Y,Z)+\eta^{i}(V)\eta^{j}(X)g(h_{i}h_{j}Z,Y)+\eta^{i}(Z)\eta^{j}(Y)g(h_{i}h_{j}V,X)$
$-\eta^{i}(Y)\eta^{j}(Z)g(h_{i}h_{j}X,V){+}\eta^{i}(Y)\eta^{j}(V)g(h_{i}h_{j}X,Z){-}\eta^{i}(V)\eta^{j}(Y)g(h_{i}h_{j}Z,X)\big]+\delta(X,Y,Z,V)=0.$
Simplification of this using the skew-symmetry of $h_{i}$ yields (46). Replacing $X$ by fX{f}X and $Y$ by fY{f}Y in (46) and using (1), we get (5).
Let $\{e_{k}\}_{k=1,\ldots,2n+1}$ be a local orthonormal frame on $M$ with $e_{2n+j}=\xi_{j}\ (1\leq j\leq s)$. Putting $X=Y=e_{k}$ in (43) for $\{e_{k}\}$, then using (12) and summing over $k=1,\ldots,2n+s$, we get
$\sum\nolimits_{i=1}^{s}\,\sum\nolimits_{k=1}^{2n+s}(\nabla_{e_{k}}h_{i})\,e_{k}=\sum\nolimits_{j=1}^{s}\sum\nolimits_{i=1}^{s}\operatorname{tr}(h_{i}h_{j})\xi_{j}.$
Next, taking the inner product of (42) with $Z$, tracing over $X$ and $Y$, and then substituting the above relation, we obtain (44). Replacing $Y$ by $h_{i}Y$ in (43), putting $Y=e_{k}$ in the gotten equation and summing over $k=1,\ldots,2n+s$, we get
$\operatorname{tr}((\nabla_{X}h_{i})h_{i})=\sum\nolimits_{k=1}^{2n+s}g((\nabla_{X}h_{i})h_{i}e_{k},e_{k})=\sum\nolimits_{j=1}^{s}\sum\nolimits_{k=1}^{2n+s}\eta^{j}(e_{k})\,g(h_{i}h_{j}X,h_{i}e_{k})=0.$
This implies $X({\rm tr}(h_{i}^{2}))=0\ (X\in\mathfrak{X}_{M})$, i.e., ${\rm tr}(h_{i}^{2})=const$. ∎
Remark 1.
The function $\delta$ of a weak nearly $\cal C$}-manifold has the following symmetries:
$\delta(Y,X,Z,V)=\delta(X,Y,V,Z)=\delta(Z,V,X,Y)=-\delta(X,Y,Z,V).$
If (13) is true, then by (45), we get
$\delta(\xi_{i},Y,Z,V)\!=\delta(X,\xi_{i},Z,V)\!=\delta(X,Y,\xi_{i},V)\!=\delta(X,Y,Z,\xi_{i})=0\quad(1\leq i\leq s).$
Lemma 6.
For a weak nearly $\cal C$}-manifold satisfying (5), (6), (9) and (13), the operators $h_{i}$ and $h_{j}$ commute for $1\leq i,j\leq s$ and each composition $h_{i}h_{j}$ is self-adjoint.
Proof.
From the first Bianchi identity, using (23), we have $R_{\xi_{i},X}\,\xi_{j}=R_{\xi_{j},X}\,\xi_{i}$ for all $X\in\mathfrak{X}_{M}$ and $1\leq i,j\leq s$. From this, applying (43) and (42), we get
$0=R_{\xi_{i},X}\,\xi_{j}-R_{\xi_{j},X}\,\xi_{i}=-(\nabla_{X}h_{i})\xi_{j}+(\nabla_{X}h_{j})\xi_{i}=h_{i}h_{j}X-h_{j}h_{i}X.$
This proves the first part, while the latter follows directly as a consequence of the commutativity and skew-symmetry of the operators $h_{i}\ (1\leq i\leq s)$. ∎
Lemma 7.
For a weak nearly $\cal C$}-manifold satisfying (5), (6), (9) and (13), we have
$g((\nabla_{X}{f})Y,{f}h_{j}Z)=\sum\nolimits_{i=1}^{s}\big[\eta^{i}(X)g(h_{j}Y,h_{i}Z)-\eta^{i}(Y)g(h_{j}X,h_{i}Z)\big]\quad(1\leq j\leq s).$ (49)
Proof.
Differentiating (19) and using $g((\nabla_{X}{f})(\nabla_{V}{f})Y,Z)$ $g((\nabla_{V}{f})Y,(\nabla_{X}{f})Z)$ gives
$g((\nabla_{V}\,{f})Y,(\nabla_{X}{f})Z)+g((\nabla_{X}{f})Y,(\nabla_{V}\,{f})Z)=g((\nabla^{2}_{V,X}\,{f}){f}Y,Z)$
$+g((\nabla^{2}_{V,X}\,{f}){f}Z,Y)-\sum\nolimits_{i=1}^{s}\big[g(h_{i}V,Z)\,g(h_{i}X,Y)+\eta^{i}(Z)\,g((\nabla_{V}h_{i})X,Y)$
$+g(h_{i}V,Y)\,g(h_{i}X,Z)+\eta^{i}(Y)\,g((\nabla_{V}h_{i})X,Z)+\nabla_{V}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)\big].$ (50)
Using (1), (25) and (31), we find $\nabla^{2}$-terms in (3):
$g((\nabla^{2}_{V,X}\,{f}){f}Z,Y)=g((\nabla^{2}_{V,{f}Z}\,{f})Y,X)=-g(R_{X,Y}V,QZ)+\sum\nolimits_{i=1}^{s}\eta^{i}(Z)\,g(R_{X,Y}V,\xi_{i})$
$\qquad-(1/2)\,g(R_{X,V}{f}Z,{f}Y)-(1/2)\,g(R_{V,Y}{f}Z,{f}X)-(1/2)\,g(R_{X,Y}{f}Z,{f}V),$ (51)
$g((\nabla^{2}_{V,X}\,{f}){f}Y,Z)=g((\nabla^{2}_{V,{f}Y}\,{f})Z,X)=-g(R_{X,Z}V,QY)+\sum\nolimits_{i=1}^{s}\eta^{i}(Y)\,g(R_{X,Z}V,\xi_{i})$
$\qquad-(1/2)\,g(R_{X,V}{f}Y,{f}Z)-(1/2)\,g(R_{V,Z}{f}Y,{f}X)-(1/2)\,g(R_{X,Z}{f}Y,{f}V).$ (52)
Substituting (3)–(3) in (3), we get
$g((\nabla_{V}\,{f})Y,(\nabla_{X}{f})Z)+g((\nabla_{X}{f})Y,(\nabla_{V}\,{f})Z)=-g(R_{X,Z}V,QY)-g(R_{X,Y}V,QZ)$
$-(1/2)\,g(R_{V,Z}{f}Y,{f}X)-(1/2)\,g(R_{X,Z}{f}Y,{f}V)-(1/2)\,g(R_{V,Y}{f}Z,{f}X)-(1/2)\,g(R_{X,Y}{f}Z,{f}V)$
$+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Y)\,g(R_{X,Z}V,\xi_{i})+\eta^{i}(Z)\,g(R_{X,Y}V,\xi_{i})-g(h_{i}V,Z)\,g(h_{i}X,Y)-\eta^{i}(Z)\,g((\nabla_{V}h_{i})X,Y)$
$-g(h_{i}V,Y)\,g(h_{i}X,Z)-\eta^{i}(Y)\,g((\nabla_{V}h_{i})X,Z)-\nabla_{V}\big(\eta(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)\big].$ (53)
Applying (46) twice followed by (43) in (3), the curvature terms involving $\xi_{i}$ vanish, and we get
$g((\nabla_{X}{f})Z,(\nabla_{V}{f})Y)+g((\nabla_{X}{f})Y,(\nabla_{V}{f})Z)+g(R_{X,Z}V,QY)+g(R_{X,Y}V,QZ)$
$-g(R_{V,Z}{f}X,{f}Y)-g(R_{X,Z}{f}V,{f}Y)+\sum\nolimits_{i=1}^{s}\big[g(h_{i}V,Y)\,g(h_{i}X,Z)+g(h_{i}V,Z)\,g(h_{i}X,Y)\big]$
$=-(1/4)\delta(X,Z,V,Y)-(1/4)\delta(V,Z,X,Y)-\sum\nolimits_{i=1}^{s}\nabla_{V}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big).$ (54)
Replacing in (3) $Z$ by fZ{f}Z and $V$ by fV{f}V, we find
$g((\nabla_{X}{f}){f}Z,(\nabla_{{f}V}\,{f})Y){+}g((\nabla_{X}{f})Y,(\nabla_{{f}V}\,{f}){f}Z){+}g(R_{X,{f}Z}{f}V,QY){-}g(R_{X,{f}Z}{f}^{2}V,{f}Y)$
$+g(R_{X,Y}{f}V,{f}QZ)-g(R_{{f}V,{f}Z}{f}X,{f}Y)+\sum\nolimits_{i=1}^{s}\big[g(h_{i}X,Y)\,g(h_{i}{f}V,{f}Z)$
$+g(h_{i}X,{f}Z)\,g(h_{i}{f}V,Y)\big]=(1/4)\,\delta(X,{f}Z,{f}V,Y)-(1/4)\,\delta({f}V,{f}Z,X,Y).$ (55)
Using (1), (19), (20), (3) and Lemma 2, we find
$g((\nabla_{X}{f}){f}Z,(\nabla_{{f}V}\,{f})Y)=g(Q(\nabla_{X}{f})Z,(\nabla_{V}\,{f})Y)-\sum\nolimits_{i,j=1}^{s}\eta^{i}(Z)\,\eta^{j}(V)g(h_{i}X,h_{j}Y)$
$\quad+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Z)g({f}h_{i}X,(\nabla_{V}\,{f})Y)-g(X,{f}h_{i}Z)\,g(V,{f}h_{i}Y)-\eta^{i}(V)\,g((\nabla_{X}{f})Z,{f}h_{i}Y)$
$\quad+g(h_{i}X,QZ)g(h_{i}V,QY)\big],$ (56)
$g((\nabla_{X}{f})Y,(\nabla_{{f}V}{f}){f}Z)=-g(Q(\nabla_{X}{f})Y,(\nabla_{V}\,{f})Z)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)g({f}h_{i}Z,(\nabla_{X}{f})Y)$
$\quad-\eta^{i}(Z)g({f}h_{i}V,(\nabla_{X}{f})Y)+g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)\big].$ (57)
From (45) and Lemma 2, we have
$g(R_{X,{f}Z}{f}V,Y)-g(R_{X,{f}Z}{f}^{2}V,{f}Y)$ $=g(R_{X,{f}Z}{f}V,Y)+g(R_{X,{f}Z}QV,{f}Y)$
$-\sum\nolimits_{i,j=1}^{s}\eta^{i}(X)\,\eta^{j}(V)\,g(h_{j}h_{i}Z,QY).$ (58)
On the other hand, from (22), (46) and (1) it follows that
$g(R_{X,Z}{f}V,{f}Y)$ $+g(R_{X,Z}{f}^{2}V,Y)+g(R_{X,{f}Z}{f}V,Y)+g(R_{{f}X,Z}{f}V,Y)=0,$ (59)
$-g(R_{{f}X,Z}{f}V,QY)$ $+\sum\nolimits_{i=1}^{s}\eta^{i}(Y)\,g(R_{{f}X,Z}{f}V,\xi_{i})=-g(R_{QX,{f}Z}V,{f}Y)$
$+\sum\nolimits_{i=1}^{s}\eta^{i}(X)\,g(R_{\xi_{i}\,,{f}Z}V,{f}Y)+(1/2)\,\delta({f}X,Z,V,{f}Y).$ (60)
Summing up the formulas (59) and (3) (and using (1), (45) and Lemma 2), we obtain
$g(R_{X,{f}Z}{f}V,Y)$ $+g(R_{QX,{f}Z}V,{f}Y)=g(R_{X,Z}QV,Y)-g(R_{X,Z}{f}V,{f}Y)+g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)$
$+\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)-\eta^{i}(Y)\,\eta^{j}(Z)\,g(h_{i}h_{j}V,QX)$
$+\eta^{i}(Z)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,X)\big]+(1/2)\,\delta({f}X,Z,V,{f}Y).$ (61)
Substituting (3) into (3), we get
$g(R_{X,{f}Z}\,{f}V,Y)-g(R_{X,{f}Z}\,{f}^{2}V,{f}Y)=g(R_{X,Z}\,QV,Y)-g(R_{X,Z}{f}V,{f}Y)$
$+\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(Z)\eta^{j}(V)\,g(h_{i}h_{j}Y,X)-\eta^{i}(Y)\eta^{j}(Z)\,g(h_{j}h_{i}X,QV)-\eta^{i}(X)\,\eta^{j}(V)\,g(h_{j}h_{i}Z,Y)\big]$
$-g(R_{\widetilde{Q}X,{f}Z}V,{f}Y)+g(R_{X,{f}Z}\widetilde{Q}V,{f}Y)+g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)+(1/2)\,\delta({f}X,Z,V,{f}Y).$ (62)
Using (45) and (46)–(5), we have
$g(R_{X,Y}{f}V,{f}Z)-g(R_{{f}V,{f}Z}{f}X,{f}Y)=g(R_{V,Z}{f}X,{f}Y)-g(R_{V,Z}QX,QY)$
$+\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}QY,Z)-\eta^{i}(X)\eta^{j}(Z)\,g(h_{i}h_{j}QY,V)$
$+\,\eta^{i}(Y)\,\eta^{j}(Z)\,g(h_{i}h_{j}QX,V)-\eta^{i}(Y)\,\eta^{j}(V)\,g(h_{i}h_{j}QX,Z)\big]$
$-\,(1/2)\,\delta({f}X,{f}Y,Z,V)-(1/2)\,\delta(X,Y,Z,V).$ (63)
Substituting (3), (3), (3) and (3) into (3), and using Lemma 2, we obtain
$g(Q(\nabla_{X}{f})Z,(\nabla_{V}{f})Y)-g(Q(\nabla_{X}{f})Y,(\nabla_{V}{f})Z)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)g((\nabla_{X}{f})Y,{f}h_{i}Z)$
$-\eta^{i}(V)\,g((\nabla_{X}{f})Z,{f}h_{i}Y)+\eta^{i}(Z)g((\nabla_{V}{f})Y,{f}h_{i}X)-\eta^{i}(Z)g((\nabla_{X}{f})Y,{f}h_{i}V)$
$+g(h_{i}X,QZ)g(h_{i}V,QY)-g(h_{i}X,Y)\,g(h_{i}V,QZ)\big]+\sum\nolimits_{i,j=1}^{s}\big[2\,\eta^{i}(Z)\eta^{j}(V)g(h_{i}h_{j}Y,X)$
$-\eta^{i}(X)\,\eta^{j}(Z)\,g(h_{i}h_{j}Y,QV)-\eta^{i}(Y)\,\eta^{j}(V)\,g(h_{i}h_{j}X,QZ)\big]+g(R_{X,Z}\,QV,Y)$
$-g(R_{X,Z}{f}V,{f}Y)+g(R_{V,Z}{f}X,{f}Y)-g(R_{V,Z}\,QX,QY)$
$=g(R_{\widetilde{Q}X,{f}Z}V,{f}Y){-}g(R_{X,{f}Z}\widetilde{Q}V,{f}Y){-}g(R_{X,{f}Z}{f}V,\widetilde{Q}Y){-}g(R_{X,Y}{f}V,{f}\widetilde{Q}Z)-g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)$
$-\sum\nolimits_{i=1}^{s}g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)-\sum\nolimits_{i,j=1}^{s}\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)-(1/2)\,\delta({f}X,Z,V,{f}Y)$
$+(1/2)\,\delta({f}X,{f}Y,Z,V)+(1/2)\,\delta(X,Y,Z,V)-(1/4)\,\delta(X,{f}Z,{f}V,Y)-(1/4)\,\delta({f}V,{f}Z,X,Y).$ (64)
Adding (3) to (3), we obtain
$2\,g((\nabla_{X}{f})Z,(\nabla_{V}{f})Y)+g(\widetilde{Q}(\nabla_{X}{f})Z,(\nabla_{V}{f})Y)-g(\widetilde{Q}(\nabla_{X}{f})Y,(\nabla_{V}{f})Z)$
$+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)g((\nabla_{X}{f})Y,{f}h_{i}Z)-\eta^{i}(V)\,g((\nabla_{X}{f})Z,{f}h_{i}Y)+\eta^{i}(Z)g((\nabla_{V}{f})Y,{f}h_{i}X)$
$-\eta^{i}(Z)g((\nabla_{X}{f})Y,{f}h_{i}V)+g(h_{i}X,QZ)g(h_{i}V,QY)+g(h_{i}V,Y)\,g(h_{i}X,Z)\big]$
$+\sum\nolimits_{i,j=1}^{s}\big[2\,\eta^{i}(Z)\eta^{j}(V)g(h_{i}h_{j}Y,X)-\eta^{i}(X)\eta^{j}(Z)g(h_{i}h_{j}Y,QV)-\eta^{i}(Y)\eta^{j}(V)g(h_{i}h_{j}X,QZ)\big]$
$+2\,g(R_{X,Z}\,V,Y)-2\,g(R_{X,Z}{f}V,{f}Y)-g(R_{V,Z}\,QX,QY)+g(R_{X,Y}V,QZ)$
$=g(R_{\widetilde{Q}X,{f}Z}V,{f}Y){-}g(R_{X,{f}Z}\widetilde{Q}V,{f}Y){-}g(R_{X,{f}Z}{f}V,\widetilde{Q}Y)-g(R_{X,Z}\,\widetilde{Q}V,Y)-g(R_{X,Z}V,\widetilde{Q}Y)$
$-g(R_{X,Y}{f}V,{f}\widetilde{Q}Z)-g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)-\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)$
$+\sum\nolimits_{i=1}^{s}\big[g(h_{i}X,Y)\,g(h_{i}V,\widetilde{Q}Z)-g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)-\nabla_{V}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)\big]$
$-(1/2)\,\delta({f}X,Z,V,{f}Y)+(1/2)\,\delta({f}X,{f}Y,Z,V)+(1/2)\,\delta(X,Y,Z,V)-(1/4)\delta(X,Z,V,Y)$
$-(1/4)\delta(V,Z,X,Y)-(1/4)\,\delta(X,{f}Z,{f}V,Y)-(1/4)\,\delta({f}V,{f}Z,X,Y).$ (65)
Swapping $X\leftrightarrow Z$ and $V\leftrightarrow Y$ in (3), then subtracting the gotten equation from (3) and using (8), we get
$\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)g((\nabla_{X}{f})Y,{f}h_{i}Z)-\eta^{i}(V)\,g((\nabla_{X}{f})Z,{f}h_{i}Y)+\eta^{i}(Z)g((\nabla_{V}\,{f})Y,{f}h_{i}X)$
$-\eta^{i}(Y)\,g((\nabla_{Z}\,{f})V,{f}h_{i}X)-\eta^{i}(Z)g((\nabla_{X}{f})Y,{f}h_{i}V)+\eta^{i}(Y)g((\nabla_{Z}{f})X,{f}h_{i}V)$
$-\eta^{i}(X)g((\nabla_{Y}\,{f})V,{f}h_{i}Z)+\eta^{i}(X)g((\nabla_{Z}{f})V,{f}h_{i}Y)\big]$
$+\sum\nolimits_{i,j=1}^{s}\big[2\,\eta^{i}(Z)\,\eta^{j}(V)g(h_{i}h_{j}Y,X)-2\,\eta^{i}(X)\,\eta^{j}(Y)g(h_{i}h_{j}V,Z)]$
$+g(R_{Z,V}\,QX,QY)+g(R_{Z,V}\,QX,Y)-g(R_{X,Y}\,QZ,V)-g(R_{X,Y}\,QZ,QV)$
$=g(R_{\widetilde{Q}X,{f}Z}V,{f}Y)-g(R_{X,{f}Z}{f}V,\widetilde{Q}Y)+g(R_{Z,V}{f}Y,{f}\widetilde{Q}X)$
$-g(R_{X,Y}{f}V,{f}\widetilde{Q}Z)-g(R_{\widetilde{Q}Z,{f}X}Y,{f}V)+g(R_{Z,{f}X}{f}Y,\widetilde{Q}V)$
$+\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(Z)\,\eta^{j}(Y)\,g(h_{i}h_{j}V,\widetilde{Q}X)-\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)\big]$
$+\sum\nolimits_{i=1}^{s}\big[g(h_{i}X,Y)\,g(h_{i}V,\widetilde{Q}Z)-g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)-g(h_{i}Z,V)\,g(h_{i}Y,\widetilde{Q}X)$
$+g({f}h_{i}Z,V)g({f}h_{i}X,\widetilde{Q}Y)\big]+(1/4)\,\delta({f}X,Z,{f}Y,V)+(1/4)\,\delta({f}X,{f}Y,Z,V)$
$+(1/2)\,\delta(X,Y,Z,V)+(1/4)\,\delta(X,{f}Z,{f}V,Y)-(1/4)\,\delta({f}Z,{f}V,X,Y)$
$+\sum\nolimits_{i,j=1}^{s}\big[\nabla_{Y}\big(\eta^{i}(X)\,g(h_{i}Z,\widetilde{Q}V)\big)-\nabla_{V}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)\big].$ (66)
Putting $\xi_{k}$ in place of $V$ in (3), then using Lemma 2, (12), $\nabla_{\xi_{k}}\big(\eta^{i}(Z)\,g(h_{i}X,\widetilde{Q}Y)\big)=0$, Remark 1 and (8), we get
$\qquad g((\nabla_{X}{f})Y,{f}h_{k}Z)=g((\nabla_{X}{f})Z,{f}h_{k}Y)+\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Z)g(h_{k}Y,h_{i}X)-\eta^{i}(Y)g(Qh_{k}Z,h_{i}X)$
$-\eta^{i}(Z)g(h_{k}\widetilde{Q}Y,h_{i}X){-}\eta^{i}(X)g(h_{i}h_{k}Y,\widetilde{Q}Z)\big]{+}g(R_{\xi_{k},Z}QX,QY){+}g(R_{\xi_{k},Z}QX,Y){-}2g(R_{\xi_{k},QZ}X,Y).$
Simplifying the above using (1) and (45), we have
$g((\nabla_{Y}{f})X,{f}h_{k}Z)+g((\nabla_{X}{f})Z,{f}h_{k}Y)=-\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Z)g(h_{k}Y,h_{i}X)$
$\quad-\eta^{i}(Y)g(Qh_{k}Z,h_{i}X)-\eta^{i}(Z)g(h_{k}\widetilde{Q}Y,h_{i}X)+2\eta^{i}(X)g(h_{i}Y,\widetilde{Q}h_{k}Z)\big].$ (67)
Swapping $X$ and $Y$ in the above equation and adding it back to (3), we have
$g((\nabla_{Y}f)Z,fh_{k}X)+g((\nabla_{X}f)Z,fh_{k}Y)=\sum\nolimits_{i=1}^{s}\big[\eta^{i}(X)g(h_{i}Y,Qh_{k}Z)+\eta^{i}(Y)g(h_{i}X,Qh_{k}Z)$
$-2\eta^{i}(Z)g(h_{i}X,h_{k}Y)+2\eta^{i}(Z)g(h_{i}X,\widetilde{Q}h_{k}Y)-2\eta^{i}(X)g(h_{i}Y,\widetilde{Q}h_{k}Z)-2\eta^{i}(Y)g(h_{i}X,\widetilde{Q}h_{k}Z)\big].$ (68)
By swapping $Z$ and $X$ in (3), and adding the resulting equation and (3), we get
$2\,g((\nabla_{Y}f)X,fh_{k}Z)=2\sum\nolimits_{i=1}^{s}\big[\eta^{i}(Y)g(h_{k}X,h_{i}Z)-\eta^{i}(X)g(h_{i}Z,h_{k}Y)\big].$
Finally, simplifying the above using (8) and the definition of $\widetilde{Q}$, we get the desired expression (49). ∎
4 Main results
In this section, we prove a splitting result (Theorem 1) for weak nearly $\cal C$}-manifolds satisfying some additional conditions such as (9) and (13). In Theorem 2, we characterize $(4+s)$-dimensional weak nearly $\cal C$}-manifolds. The consequences of Theorems 1 and 2 (without the requirement of (9) and (13)) present new results for nearly $\cal C$}-manifolds.
The following result generalizes [18, Proposition 4.1] (and [14, Proposition 4.2]) to the case $s>1$.
Proposition 1.
For a weak nearly $\cal C$}-manifold with conditions (5), (6), (9) and (13), the eigenvalues (and their multiplicities) of the symmetric operators $h_{i}h_{j}$ and $h_{i}^{2}$ are constant for all $1\leq i\neq j\leq s$. Moreover, each $h_{i}h_{j}$ and $h_{i}^{2}$ shares the same eigen-frame.
Proof.
From (45) and Lemma 5 we obtain
$(\nabla_{X}\,h_{i}h_{j})Y=h_{i}(\nabla_{X}\,h_{j})Y+(\nabla_{X}\,h_{i})h_{j}Y=\sum\nolimits_{k=1}^{s}\big[g(X,h_{k}h_{i}h_{j}Y)\,\xi_{k}-\eta^{j}(Y)\,h_{i}h_{j}h_{k}X\big].$ (69)
Consider an eigenvalue $\mu$ of $h_{i}h_{j}$ and a local unit vector field $Y\in\mathcal{D}$ such that $h_{i}h_{j}Y=\mu Y$. Applying (69) for any nonzero vector fields $X,Y\in\mathcal{D}$, we find $g((\nabla_{X}\,h_{i}h_{j})Y,Y)=0$, thus
$0=g((\nabla_{X}\,h_{i}h_{j})Y,Y)=g(\nabla_{X}\,(h_{i}h_{j}Y),Y)-g(h_{i}h_{j}(\nabla_{X}\,Y),Y)$
$\ \ =X(\mu)\,g(Y,Y)+\mu\,g(\nabla_{X}\,Y,Y)-g(\nabla_{X}\,Y,h_{i}h_{j}Y)=X(\mu)\,g(Y,Y),$
which implies that $X(\mu)=0$ for all $X\in\mathfrak{X}_{M}$. When $j=i$, we get a similar claim for $h_{i}^{2}$.
Next, by Lemma 6, the operators $h_{i}$ and $h_{j}$ commute for all $i,j$, which implies that $h_{i}^{2}$ and $h_{j}^{2}$ also commute. Each $h_{i}^{2}$ is self-adjoint, and therefore, diagonalizable.
It is well known that two diagonalizable operators on a finite-dimensional Euclidean space are simultaneously diagonalizable if and only if they commute. This fact extends by mathematical induction to any finite family of commuting diagonalizable operators. By the above, there exists an orthonormal basis $\{e_{k}\}$ of the tangent space $TM$ such that $h_{i}^{2}e_{k}=\lambda_{i,k}\,e_{k}$ for all $1\leq i\leq s$ and $k$. ∎
By Proposition 1, the spectrum of the self-adjoint operator $h_{i}^{2}$ has the form
$Spec(h_{i}^{2})=\{0,-\lambda_{i,1}^{2},\ldots-\lambda_{i,r_{i}}^{2}\}\qquad(1\leq i\leq s),$ (70)
where $\lambda_{i,j}$ is a positive real number and $\lambda_{i,j}\neq\lambda_{i,k}$ for $j\neq k$. If $X\neq 0$ is an eigenvector of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$, then $X,{f}X,h_{i}X$ and $h_{i}\,{f}X$ are orthogonal nonzero eigenvectors of $h_{i}^{2}$ with the same eigenvalue $\lambda^{2}_{i,j}$. Since $h_{i}(\xi_{j})=0$ for all $1\leq i,j\leq s$, the Reeb distribution $\ker f=\operatorname{span}\{\xi_{1},\dots,\xi_{s}\}$ is contained in the kernel of $h_{i}$. Consequently, the eigenvalue 0 of $h_{i}^{2}$ has multiplicity $2p_{i}+s$ for some integer $p_{i}\geq 0$, associated with each operator $h_{i}^{2}$. Denote by
$\bullet$ $D_{i,0}$ the smooth distribution of the eigenvectors of $h_{i}^{2}$ with eigenvalue 0 orthogonal to $\ker f$;
$\bullet$ $D_{i,j}\ (1\leq j\leq r_{i})$ be the smooth distribution of the eigenvectors of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$.
Note that $D_{i,0}$ and $D_{i,j}$ belong to $\cal D$} and are f{f}- and $h_{k}$-invariant for each $1\leq i,k\leq s$.
Remark 2.
By the skew-symmetry of $h_{i}$, for every $X,Y\in\mathfrak{X}_{M}$, we have
$d\eta^{i}(X,Y)$ $=g(\nabla_{X}Y,\xi_{i})+g(Y,\nabla_{X}\xi_{i})-g(\nabla_{Y}X,\xi_{i})-g(X,\nabla_{Y}\xi_{i})-g([X,Y],\xi_{i})$
$=g(Y,h_{i}X)-g(X,h_{i}Y)=2\,g(h_{i}X,Y)\quad(1\leq i\leq s).$ (71)
The above shows that $\operatorname{ker}h_{i}=\operatorname{ker}d\eta^{i}$ and as a consequence of this, $\operatorname{dim\,(ker}d\eta^{i})\geq s$ for all $i$.
The following result generalizes [18, Proposition 4.3] (and [14, Proposition 4.1]) to the case $s>1$.
Proposition 2.
[…]
Proof.
By (2) and the condition ${\bf h}\equiv 0$, we have $d\eta^{i}=0$ for all $i$. We have $0=d\eta^{i}(X,Y)=-\eta^{i}([X,Y])$ for all $i$ and $X,Y\in{\cal D}$, which indicates that $X,Y]\in{\cal D}$; thus the distribution $\cal D$} is integrable. Integral manifolds of $\cal D$} are totally geodesic as $g(\nabla_{X}\,Y,\xi_{i})=-g(Y,h_{i}X)=0$ for $X,Y\in{\cal D}$ and $1\leq i\leq s$. Since $\operatorname{ker}f$ is integrable, totally geodesic and flat (see Lemma 1), by de Rham Decomposition Theorem, e.g., [12], the manifold is locally the Riemannian product $\mathbb{R}^{s}\times\bar{M}$. The weak nearly $\cal C$}-structure on $M$ induces on $\bar{M}$ a weak nearly Kähler structure.
Conversely, if a weak nearly $\cal C$}-manifold is locally the Riemannian product $\mathbb{R}^{s}\times\bar{M}$, where $\bar{M}$ is a weak nearly Kähler manifold and $\xi_{i}=(\partial_{x_{i}},0)$ ($(x,y)\in\mathbb{R}^{s}\times\bar{M}$), then $d\eta^{i}=0$ for $1\leq i\leq s$ or simply ${\bf h}\equiv 0$, by (2). By this and the definition of $\xi_{i}$, the conditions (5) and (6) are true. ∎
The following result generalizes [18, Proposition 4.2] (and [14, Proposition 4.3]) to the case $s>1$.
Proposition 3.
Let $(M^{\,2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ be a weak nearly $\mathcal{C}$-manifold with conditions (5), (6), (9) and (13). Then
  • (i)
    distributions $\ker f\oplus{\cal D}_{i,j}\ (1\leq i\leq s,\,1\leq j\leq r_{i})$ are integrable with totally geodesic leaves.
Moreover, if $\dim(\ker d\eta^{i})>s$ for all $i$, then the following is true:
  • (ii)
    ${\cal D}_{i,0}\neq 0$ and distributions $\operatorname{ker}f\oplus{\cal D}_{i,0}$ are integrable with totally geodesic leaves;
  • (iii)
    if the exterior derivatives of the coframe $\{\eta^{i}\}$ have a common kernel, then distributions ${\cal D}_{i,0}$ coincide, are integrable with totally geodesic leaves, and the induced structure on each leaf of ${\cal D}_{i,0}$ is a weak nearly Kähler structure $(\bar{f},\bar{g})$ with the property $\bar{\nabla}(\bar{f}^{\,2})=0$;
  • (iv)
    the distribution $\ker f\oplus{\cal D}_{i,1}\oplus\ldots\oplus{\cal D}_{i,r_{i}}$ is integrable with totally geodesic leaves.
Proof.
Consider an eigenvector $X$ of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$, see (70). Then $\nabla_{X}\,\xi_{k}=h_{k}X\in{\cal D}_{i,j}$ ($k=1,\ldots,s$), as $h_{i}^{2}(h_{k}X)=h_{k}(h_{i}^{2}X)=-\lambda^{2}_{i,j}\,h_{k}X$. Since (69) implies $\nabla_{\xi_{k}}h_{i}^{2}=0$, then $\nabla_{\xi_{k}}X$ is also an eigenvector of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$ for all $k=1,..,s$. By Lemma 1, the distribution $\operatorname{ker}f$ defines a flat totally geodesic foliation. Moreover, since $\nabla_{X}\,\xi_{k},\ \nabla_{\xi_{k}}X\in\mathcal{D}_{i,j}$ for all $X\in\mathcal{D}_{i,j}$, we have $\operatorname{ker}f\oplus\mathcal{D}_{i,j}$ is integrable (involutive). Now, taking $X,Y\in{\cal D}_{i,j}$ and applying (69), we get
$h_{i}^{2}(\nabla_{X}\,Y)=-\lambda^{2}_{i,j}\,\nabla_{X}Y-(\nabla_{X}\,h_{i}^{2})Y=-\lambda^{2}_{i,j}\,\nabla_{X}Y+\lambda^{2}_{i,j}\sum\nolimits_{k=1}^{s}g(X,h_{i}Y)\,\xi_{k}.$
Using the above and the anti-commutativity of $f$ and $h_{i}$, we have
$h_{i}^{2}({f}^{2}\nabla_{X}Y)={f}^{2}(h_{i}^{2}\nabla_{X}Y)=-\lambda^{2}_{i,j}\,{f}^{2}(\nabla_{X}Y);$
hence, ${f}^{2}\nabla_{X}Y\in{\cal D}_{i,j}$. Similarly, using (18), we get $\widetilde{Q}\nabla_{X}Y\in{\cal D}_{i,j}$. It follows that
$\nabla_{X}Y=-\widetilde{Q}\,\nabla_{X}Y-{f}^{2}\nabla_{X}Y+\sum\nolimits_{k=1}^{s}\eta^{k}(\nabla_{X}Y)\,\xi_{k},$
see (1), belongs to the distribution $\ker f\oplus{\cal D}_{i,j}$. This proves $(i)$.
[…]
Since $\mathcal{D}_{i,0}\neq 0$, we get $h_{i}^{\,2}(\nabla_{X}Y)=0$ for all $X,Y\in\mathcal{D}_{i,0}$. In addition, we compute $g(\nabla_{X}Y,\xi_{k})=-g\big(Y,\nabla_{X}\xi_{k}\big)=-g(Y,h_{k}X)$, which vanishes if $\operatorname{ker}h_{i}=\operatorname{ker}h_{k}$ for all $1\leq i,k\leq s$. By (2), this condition is equivalent to the fact that the exterior derivatives of the coframe $\{\eta^{i}\}$ have a common kernel. Thus, each distribution $D_{i,0}$ defines a totally geodesic foliation if and only if the exterior derivatives of the coframe $\{\eta^{i}\}$ have a common kernel. By (8) and (18), the leaves of $D_{i,0}$ are f{f}-invariant and $Q$-invariant. Thus, the weak nearly $\cal C$}-structure on $M$ with conditions (5), (6), (9) and (13) induces a weak nearly Kähler structure $(\bar{f},\bar{g})$ on each leaf of $D_{i,0}$ with the property $\bar{\nabla}(\bar{f}^{\,2})=0$, where $\bar{\nabla}$ is the Levi-Civita connection of $\bar{g}$. This proves $(iii)$.
To establish $(iv)$, it suffices to verify that for $X\in{\cal D}_{i,j}$ and $Y\in{\cal D}_{i,k}$ with $j\neq k$, one has $\nabla_{X}Y\in\ker f\oplus{\cal D}_{i,1}\oplus\cdots\oplus{\cal D}_{i,r}$. Similarly to part $(i)$, we obtain $h_{i}^{2}(\nabla_{X}Y)=-\lambda_{i,k}\,\nabla_{X}Y+\lambda_{i,k}\sum_{j=1}^{s}g(X,h_{j}Y)\,\xi_{j}$; hence, $h_{i}^{2}(f^{2}\nabla_{X}Y)=-\lambda_{i,k}\,f^{2}\nabla_{X}Y$. Thus $f^{2}\nabla_{X}Y\in{\cal D}_{i,k}$, and we calculate $\widetilde{Q}\nabla_{X}Y\in{\cal D}_{i,k}$. Combining these observations, we may write
$\nabla_{X}Y=-\widetilde{Q}(\nabla_{X}Y)-f^{2}\nabla_{X}Y+\sum\nolimits_{k=1}^{s}\eta^{k}(\nabla_{X}Y)\,\xi_{k}\in\operatorname{ker}f\oplus{\cal D}_{i,k}\subset\ker f\oplus{\cal D}_{i,1}\oplus\cdots\oplus{\cal D}_{i,r}.$
Similarly, we obtain $\nabla_{Y}X\in\operatorname{ker}f\oplus{\cal D}_{i,j}$. Consequently, both $[X,Y]$ and $\nabla_{X}Y$ belong to $\ker f\oplus{\cal D}_{i,1}\oplus\cdots\oplus{\cal D}_{i,r}$, which completes the proof. ∎
Definition 4 (see [9]).
Let $M$ be a $(2n+s)$-dimensional smooth manifold where $n,s$ are positive integers. An s-contact structure on $M$ is a collection $\vec{\eta}=(\eta^{1},\cdots,\eta^{s})$ of $s>1$ pointwise linearly independent non-vanishing one-forms, together with a splitting $TM={\mathcal{R}}\oplus{\cal D}$ of the tangent bundle, satisfying the following conditions:
  • (i)
    ${\cal D}:=\bigcap_{\,i}\ker\eta^{i}$;
  • (ii)
    $d\eta^{i}|_{\cal D}$ is non-degenerate for every $i$;
  • (iii)
    $\ker d\eta^{i}={\mathcal{R}}$ for every $i$.
A manifold $M$ endowed with such a structure is called an $s$-contact manifold and is denoted by $(M,\vec{\eta},{\mathcal{R}}\oplus{\cal D})$. A coframe $\vec{\eta}$ (and the $s$-contact structure it defines) is said to be uniform if it satisfies
  • (iv)
    $d\eta^{i}=d\eta^{j}$ for all $1\leq i,j\leq s$.
An example of $s$-contact manifolds is a weak $f$-K-contact manifold, in fact, it is uniform. By (2), $d\eta^{i}=d\eta^{j}\iff h_{i}=h_{j}$ for all $1\leq i,j\leq s$.
Next, we will use Lemmas 37 to characterize $(4+s)$-dimensional weak nearly $\cal C$}-manifolds. But first, we generalize Proposition 3.2 in [14] to the case $s>1$.
Proposition 4.
Let $\Omega^{k}(M)$ be the space of $k$-forms on an $s$-contact manifold $(M^{2n+s},\vec{\eta},{\mathcal{R}}\oplus{\cal D})$. If $n\geq 3$, then the map $\mathcal{E}_{d\eta^{i}}:\ \Omega^{2}(M)\longrightarrow\Omega^{4}(M),\,\beta\mapsto d\eta^{i}\wedge\beta$, is injective for all $i=1,\ldots,s$.
Proof.
We denote by $\Omega^{k}(\mathcal{D})$ the space of $k$-forms along $\mathcal{D}$, i.e., $k$-forms vanishing whenever at least one argument lies in $\mathcal{R}$}. The proof of the proposition consists of four steps.
Step 1 (Symplecticity on $\mathcal{D}$ and Lefschetz injectivity): By the definition of an $s$-contact manifold, $d\eta^{i}|_{\mathcal{D}}$ is nondegenerate at every point $p\in M$. Each fibre $(\mathcal{D}_{p},d\eta^{i}|_{{\mathcal{D}}_{p}})$ is a $2n$-dimensional symplectic vector space, hence $(d\eta^{i}|_{{\mathcal{D}}})^{n}$ is nowhere vanishing. In particular, by [6, Proposition 1.1] (the algebraic Lefschetz property), the linear maps $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\,\Omega^{2}(\mathcal{D})\longrightarrow\Omega^{4}(\mathcal{D})$, $\mathcal{E}_{d\eta^{i}|_{{\mathcal{D}}}}:\,\Omega^{1}(\mathcal{D})\longrightarrow\Omega^{3}(\mathcal{D})$ are injective for $n\geq 3$ and $n\geq 2$, respectively (the latter injectivity, in particular, holds when $n\geq 3$).
Step 2 (Bundle-level decomposition by $\eta$-degree): At each $p\in M$ we have a direct-sum decomposition $T_{p}^{*}M={\mathrm{span}\{\eta^{1}_{p},\dots,\eta^{s}_{p}\}}\ \oplus\ \mathcal{D}_{p}^{*},$ which induces a decomposition of exterior powers: $\Lambda^{k}(T_{p}^{*}M)\;=\;\bigoplus\nolimits_{j=0}^{\min(k,s)}\Lambda^{j}(\mathrm{span}\{\eta^{1}_{p},\dots,\eta^{s}_{p}\})\,\wedge\,\Lambda^{k-j}\mathcal{D}_{p}^{*}$. Since the ranks are constant and the summands depend smoothly on $p$, this yields a decomposition of bundles and spaces of sections:
$\Omega^{k}(M)\;=\;\bigoplus\nolimits_{j=0}^{\min(k,s)}{\Lambda^{j}\!\big(\mathrm{span}\{\eta^{1},\dots,\eta^{s}\}\big)}\wedge\Omega^{k-j}(\mathcal{D}).$
In particular, every $\beta\in\Omega^{2}(M)$ admits a unique splitting
$\beta\;=\;\beta_{0}\;+\;\sum\nolimits_{a=1}^{s}\eta^{a}\wedge\mu_{a}\;+\;\sum\nolimits_{1\leq a<b\leq s}f_{ab}\,\eta^{a}\wedge\eta^{b},$ (72)
with $\beta_{0}\in\Omega^{2}(\mathcal{D})$, $\mu_{a}\in\Omega^{1}(\mathcal{D})$ and $f_{ab}\in C^{\infty}(M)$.
Step 3 ($d\eta^{i}$ has no $\eta$-components): Expand $d\eta^{i}$ according to the same decomposition:
$d\eta^{i}\;=\;d\eta^{i}|_{\mathcal{D}}+\sum\nolimits_{a=1}^{s}\eta^{a}\wedge\gamma_{a}\;+\;\sum\nolimits_{1\leq a<b\leq s}h_{ab}\,\eta^{a}\wedge\eta^{b},$
with $\gamma_{a}\in\Omega^{1}(\mathcal{D})$ and $h_{ab}\in C^{\infty}(M)$. Let $\{\xi_{1},\dots,\xi_{s}\}$ be a local frame of $\mathcal{R}$} dual to $\{\eta^{1},\dots,\eta^{s}\}$. Using $\ker\,d\eta^{i}={\mathcal{R}}$, we have $\iota_{\xi_{a}}d\eta^{i}=0$ for each $i$ and $a$. Contracting the above expansion yields
$0=\iota_{\xi_{a}}d\eta^{i}=\ \gamma_{a}\;+\;\sum\nolimits_{b}h_{ab}\,\eta^{b}\;-\;\sum\nolimits_{b}h_{ba}\,\eta^{b}=\ \gamma_{a}\;+\;2\sum\nolimits_{b}h_{ab}\,\eta^{b}.$
Since $\{\eta^{b}\}$ are pointwise independent, it follows that $\gamma_{a}=0$ and $h_{ab}=0$ for all $a,b$. Hence $d\eta^{i}=d\eta^{i}|_{\mathcal{D}}$, that means that $d\eta^{i}$ has no terms involving $\eta^{i}$.
Step 4 (Wedge and compare $\eta$-degrees): Assume $d\eta^{i}\wedge\beta=0$. Using (72),
$0=d\eta^{i}|_{\mathcal{D}}\wedge\beta$ $={d\eta^{i}|_{\mathcal{D}}\wedge\beta_{0}}+\sum\nolimits_{a=1}^{s}{\eta^{a}\wedge\bigl(d\eta^{i}|_{\mathcal{D}}\wedge\mu_{a}\bigr)}\;+\;\sum\nolimits_{1\leq a<b\leq s}{f_{ab}\,\eta^{a}\wedge\eta^{b}\wedge d\eta^{i}|_{\mathcal{D}}}.$ (73)
Since different $\eta$-degrees lie in a direct sum, each component vanishes separately:
$\bullet$ Zero $\eta$-part (first term of (73)): $d\eta^{i}|_{\mathcal{D}}\wedge\beta_{0}=0$ in $\Omega^{4}({\mathcal{D}})$. Now, as $d\eta^{i}|_{\mathcal{D}}\in\Omega^{2}({\mathcal{D}})$ and by the decomposition (72), $\beta_{0}\in\Omega^{2}({\mathcal{D}})$ and since $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\Omega^{2}({\mathcal{D}})\to\Omega^{4}({\mathcal{D}})$ is injective for $n\geq 3$ (by [6, Proposition 1.1]), we get $\beta_{0}=0$.
$\bullet$ One $\eta$-part: For each $a$, $d\eta^{i}|_{\mathcal{D}}\wedge\mu_{a}=0$ in $\Omega^{3}({\mathcal{D}})$. Again, since $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\Omega^{1}({\mathcal{D}})\to\Omega^{3}({\mathcal{D}})$ is injective (in particular, for $n\geq 3$), $\mu_{a}=0$ for all $a$.
$\bullet$ Two $\eta$-parts: From Step 4 we are left with $\sum\nolimits_{1\leq a<b\leq s}f_{ab}\,\eta^{a}\wedge\eta^{b}\wedge d\eta^{i}|_{\mathcal{D}}=0$. Wedge this identity with $(d\eta^{i}|_{\mathcal{D}})^{\,n-1}$ to obtain $\sum\nolimits_{a<b}f_{ab}\,\eta^{a}\wedge\eta^{b}\wedge(d\eta^{i}|_{\mathcal{D}})^{\,n}=0$. By Step 1, $(d\eta^{i}|_{\mathcal{D}})^{\,n}$ is a nowhere-vanishing horizontal top form on $\mathcal{D}$}. Fix any point $p\in M$. Then $\{\,\eta^{a}_{p}\wedge\eta^{b}_{p}\,\}_{a<b}$ is a basis of $\Lambda^{2}(\mathrm{span}\{\eta^{1}_{p},\dots,\eta^{s}_{p}\})$, and wedging with $(d\eta^{i}|_{\mathcal{D}})^{\,n}\neq 0$ multiplies these basis elements by the same nonzero horizontal volume form. Hence, the linear independence is preserved, and the equality above forces $f_{ab}(p)=0$ for $a<b$. Since $p\in M$ is arbitrary, $f_{ab}\equiv 0$ on $M$. Thus, $\beta=\beta_{0}+\sum_{a}\eta^{a}\wedge\mu_{a}+\sum_{a<b}f_{ab}\,\eta^{a}\wedge\eta^{b}=0$, and $\mathcal{E}_{d\eta^{i}}$ is injective on $\Omega^{2}(M)$ for $n\geq 3$. ∎
The following result generalizes [18, Theorem 4.2] (see also Theorem 4.4 in [14]) to the case $s>1$.
Theorem 1.
Let $(M^{\,2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ be a weak nearly $\cal C$-manifold satisfying (5), (6), (9) and (13). If the coframe $\{\eta^{i}\}$ is uniform (has a common exterior derivative), then $n=2$ and $M$ admits a weak $f$-K-contact structure $(\hat{f},\hat{Q},\vec{\xi},\vec{\eta},g)$ defined by $\hat{f}=-\nabla\xi_{i}$ and $\hat{Q}X=R_{\xi_{i}}X\ (X\in{\cal D})$.
Proof.
We consider 2-forms $\Phi^{(k)}_{i}(X,Y)=g({f}h_{i}^{k}X,Y)$ and $\Phi^{(1)}_{i,j}(X,Y)=g({f}h_{i}h_{j}X,Y)$, where $i=1,\dots s,\ k=0,1$. Also, since $h_{i}$ and $h_{j}$ commute, we have $\Phi^{(1)}_{i,j}=\Phi^{(1)}_{j,i}$. It is easy to calculate $d\Phi(X,Y,Z)=g((\nabla_{X}{f})Z,Y)+g((\nabla_{Y}\,{f})X,Z)+g((\nabla_{Z}\,{f})Y,X)$, see [15]. We will show that
$d\Phi^{(0)}_{i}=3\sum\nolimits_{j=1}^{s}\eta^{j}\wedge\Phi^{(1)}_{j},\quad d\Phi^{(1)}_{i}=3\sum\nolimits_{j=1}^{s}\eta^{j}\wedge\Phi^{(1)}_{i,j}$. (74)
Indeed, applying (14) and ${f}\,\xi_{i}=0$, we find the $\xi_{i}$-component of $(\nabla_{X}{f})Y$:
$g((\nabla_{X}{f})Y,\xi_{i})=-g((\nabla_{X}{f})\,\xi_{i},Y)=g({f}\nabla_{X}\,\xi_{i},Y)=g({f}h_{i}X,Y).$ (75)
Replacing $Z$ by fZ{f}Z in (49) and using (8), we obtain for all $1\leq i\leq s$
$-g((\nabla_{X}\,{f})Y,{f}^{2}h_{i}Z)=\sum\nolimits_{j=1}^{s}\big[\eta^{j}(X)\,g(h_{i}fZ,h_{j}Y)-\eta^{j}(Y)\,g(h_{j}X,{f}h_{j}Z)\big].$ (76)
By conditions, ${\bf h}\not\equiv 0$ on ${\cal D}\setminus\{0\}$, thus from (76) we get $g((\nabla_{X}\,{f})Y,V)=0$ for $X,Y,V\in{\cal D}$. By the above and (75), using $X=X^{\top}+\sum\nolimits_{i=1}^{s}\eta^{i}(X)\,\xi_{i}$ and $Y=Y^{\top}+\sum\nolimits_{i=1}^{s}\eta^{i}(Y)\,\xi_{i}$, we obtain
$g((\nabla_{X}\,{f})Y,V)$
$=\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)\,g((\nabla_{X^{\top}}\,{f})Y^{\top},\xi_{i})+\eta^{i}(X)\,g((\nabla_{\,\xi_{i}}\,{f})Y^{\top},V)+\eta^{i}(Y)\,g((\nabla_{X^{\top}}\,{f})\,\xi_{i},V)\big]$
$=\sum\nolimits_{i=1}^{s}\big[-\eta^{i}(V)\,g((\nabla_{X^{\top}}\,{f})\,\xi_{i},Y^{\top})-\eta^{i}(X)\,g((\nabla_{Y^{\top}}\,{f})\,\xi_{i},V)+\eta^{i}(Y)\,g((\nabla_{X^{\top}}\,{f})\,\xi_{i},V)\big]$
$=\sum\nolimits_{i=1}^{s}\big[\eta^{i}(V)\,g({f}h_{i}X,Y)+\eta^{i}(X)\,g({f}h_{i}Y,V)+\eta^{i}(Y)\,g({f}h_{i}V,X)\big].$ (77)
Using the expression for $d\Phi_{i}^{(0)}$ and (4), we have the first equality of (74).
Similarly, using (4) and (45), we get
$g((\nabla_{X}\,({f}\,h_{i}))Y,Z)=g((\nabla_{X}{f})h_{i}Y,Z)+g({f}(\nabla_{X}\,h_{i})Y,Z)$
$=\sum\nolimits_{j=1}^{s}\big[\eta^{j}(X)\,g({f}\,h_{i}h_{j}Y,Z)+\eta^{j}(Y)\,g({f}\,h_{i}h_{j}Z,X)+\eta^{j}(Z)\,g({f}\,h_{i}h_{j}X,Y)\big],$
which implies the second equality of (74). From (74), using the symmetry: $\Phi^{(1)}_{i,j}=\Phi^{(1)}_{j,i}$ and the anti-symmetry of wedge products: $\eta^{j}\wedge\eta^{i}=-\eta^{i}\wedge\eta^{j}$, we obtain
$0=d^{2}\Phi^{(0)}_{i}=3\sum\nolimits_{j=1}^{s}\big[d\eta^{j}\wedge\Phi^{(1)}_{j}-\eta^{j}\wedge d\Phi^{(1)}_{j}\big]=3\sum\nolimits_{j=1}^{s}d\eta^{j}\wedge\Phi^{(1)}_{j}$. (78)
Substituting $Y=\xi_{i}$ into (45) and applying (15), we represent the $\xi$-sectional curvature as
$K(\xi_{i},X)=g(h_{i}X,h_{i}X)=\|h_{i}X\|^{2}\quad(X\in\mathcal{D},\ g(X,X)=1,\ 1\leq i\leq s).$
Under the assumptions, the $\xi$-sectional curvature of $(M,g)$ is positive, which implies that the Jacobi operators $R_{\,{\xi_{i}}}:X\mapsto R_{X,\,{{\xi_{i}}}}\,{{\xi_{i}}}\ (1\leq i\leq s)$ are positive definite on $\mathcal{D}$. By [17, Theorem 2], there exists a weak $f$-K-contact structure $(\hat{f},\hat{Q},\vec{\xi},\vec{\eta},g)$ on $M$ defined by $\hat{f}=-\nabla\xi_{i}$ and $\hat{Q}X=R_{\xi_{i}}X\ (X\in{\cal D})$.
As a consequence of this, $\eta^{i}$ is a contact form and $d\eta^{i}$ is non-degenerate on $\mathcal{D}$. Thus, $M$ is a uniform $s$-contact manifold. Now, if $\dim M>4+s$, then (since $M$ is a uniform $s$-contact manifold) (78) implies that $d\eta^{j}\wedge\sum_{j=1}^{s}\Phi^{(1)}_{j}=0\,\text{for all }j.$ By Proposition 4, this condition forces $\Phi^{(1)}_{i}=0$ for all $i$; hence, ${\bf h}\equiv 0$. This contradicts the fact that $d\eta^{i}$ is non-degenerate on $\mathcal{D}$. Therefore, $\dim M\leq 4+s$. Finally, since $d\eta^{i}$ is non-degenerate on $\mathcal{D}$, the distribution $\operatorname{ker}d\eta^{i}$ has dimension $s$. We conclude that $h_{i}^{2}|_{\cal D}$ is non-zero and has at least one non-zero eigenvalue, say, $\lambda^{2}$, whose multiplicity is $4$: if $X$ is an eigenvector of $h_{i}^{2}$ with eigenvalue $\lambda^{2}$, then $\{X,fX,h_{i}X,h_{i}fX\}$ are also orthogonal eigenvectors of $h_{i}^{2}$ with the same eigenvalue. Therefore, $M$ is a $(4+s)$-dimensional manifold and the multiplicity of the eigenvalue $\lambda^{2}$ of $h_{i}^{2}$ is $4$. ∎
Remark 3.
It should be noted that the weak $f$-K-contact structure in Theorem 1 does not coincide with the existing weak nearly $\mathcal{C}$-structure $({f},Q,\vec{\xi},\vec{\eta},g)$, even though the $1$-forms $\eta^{i}$ are identical. Nevertheless, each $\eta^{i}$ is a contact form, thus $d\eta^{i}$ is non-degenerate on $\mathcal{D}$.
Corollary 2.
Let a nearly $\cal C$-manifold $M^{\,2n+s}({f},\vec{\xi},\vec{\eta},g)$ satisfy (5) and (6). If the coframe $\{\eta^{i}\}$ is uniform, then $n=2$ and $M$ admits an $f$-K-contact structure $(\hat{f},\vec{\xi},\vec{\eta},g)$ with $\hat{f}=-\nabla\xi_{i}$.
We generalize the splitting Theorem 4.1 in [18] (see also Theorem 4.5 in [14]) to the case $s>1$.
Theorem 2.
[…]
$(i)~\mathbb{R}^{s}\times\bar{M}^{\,2n},\quad(ii)~B^{\,4+s}\times\bar{M}^{\,2n-4},$ (79)
[…]
Proof.
If ${\bf h}\equiv 0$, then by Proposition 2, $M$ is locally isometric to $\mathbb{R}^{s}\times\bar{M}^{\,2n}$.
[…]
Corollary 3.
Let $(M^{\,2n+s},{f},\vec{\xi},\vec{\eta},g)$ be a nearly $\cal C$}- (non-$\cal C$}-) manifold of $n>2$ with conditions (5) and (6). If the coframe $\{\eta^{i}\}$ is uniform, then $M$ is locally isometric (the isometry is global if $M$ is complete simply connected) to one of the Riemannian products $\mathbb{R}^{s}\times\bar{M}^{\,2n}$ and $B^{\,4+s}\times\bar{M}^{\,2n-4}$, where the induced structure on $\bar{M}$ is a nearly Kähler structure, and the induced structure on $B$ is a nearly $\cal C$}- (non-$\cal C$}-) structure.
Example 3.
Let’s illustrate (79)$(ii)$ of Theorem 2 in the case of $s=1$ (for (79)$(i)$ with $s>1$ see Example 1(iii)(iii)). Let $(f_{B},\xi_{B},\eta_{B},g_{B})$ be the nearly cosymplectic (not cosymplectic, since $d\eta_{B}\neq 0$) structure on $B^{5}=\mathbb{S}^{5}$ obtained as a totally geodesic hypersurface of the nearly Kähler six-dimensional sphere, see [4]. Define $F_{B}:=\lambda\,f_{B}$, where $\lambda\,\in\mathbb{R}_{>0}$, and $\tilde{Q}:=\lambda^{2}\,\mathrm{Id}_{B}+(1-\lambda^{2})\,\eta_{B}\otimes\xi_{B}$. Then $(F_{B},\tilde{Q},\xi_{B},\eta_{B},g_{B})$ is a weak nearly cosymplectic structure as $(\nabla^{B}_{X}F_{B})X=\lambda\,(\nabla^{B}_{X}f_{B})X=0$ for all $X\in\mathfrak{X}_{B}$, where $\nabla^{B}$ is the Levi-Civita connection on $B$. Note that, since $d\eta_{B}\neq 0$, $(F_{B},\tilde{Q},\xi_{B},\eta_{B},g_{B})$ is not weak cosymplectic. Let $(N^{2n-4},J_{N},g_{N})$ be a weak nearly Kähler manifold satisfying $J_{N}^{2}=-\lambda^{2}\operatorname{Id}_{TN}$ with $\lambda\in\mathbb{R}_{>0}$. Let $\nabla^{N}$ be the Levi-Civita connection on $N$. We equip $M=B^{5}\times N^{2n-4}$ with the product metric $g=g_{B}\oplus g_{N}$. Then the tangent bundle splits orthogonally as $TM=TB\oplus TN$, and the Levi-Civita connection on $M$ satisfies $\nabla_{X}Y=\big(\nabla^{B}_{X_{B}}Y_{B},\,\nabla^{N}_{X_{N}}Y_{N}\big)$ for $X=(X_{B},X_{N})$ and $Y=(Y_{B},Y_{N})$. The mixed derivatives on $(M,g)$ disappear since the factors are totally geodesic. For any $X=X_{B}+X_{N}\in\mathfrak{X}_{M}$, we define
$\xi:=\xi_{B},\quad\eta(X):=\eta_{B}(X_{B}),\quad f(X):=F_{B}X_{B}+J_{N}X_{N},\quad Q:=\lambda^{2}\,\mathrm{Id}+(1-\lambda^{2})\eta\otimes\xi.$
Thus $f$ acts as $F_{B}$ on the $B^{5}$-factor and as $J_{N}$ on the $N^{2n-4}$-factor; $\xi$ and $\eta$ are lifted from the $B^{5}$-factor; $\eta$ annihilates $TN^{2n-4}$: $\eta\,|_{TN}=0.$ For $X=X_{B}+X_{N}$, we have
$f^{2}(X)=F_{B}^{2}X_{B}+J_{N}^{2}X_{N}=-\lambda^{2}X_{B}-(1-\lambda^{2})\eta_{B}(X_{B})\xi_{B}+\eta_{B}(X_{B})\xi_{B}-\lambda^{2}X_{N}=-QX+\eta(X)\xi.$
Hence $f^{2}=-Q+\eta\otimes\xi$ is true, where $Q$ is a self-adjoint (1,1)-tensor. Moreover, since $g_{B}$ and $g_{N}$ satisfy their respective Hermitian identities, $g_{B}(F_{B}X_{B},F_{B}Y_{B})=g_{B}(QX_{B},Y_{B})-\eta_{B}(X_{B})\eta_{B}(Y_{B})$ and $g_{N}(J_{N}X_{N},J_{N}Y_{N})=\lambda^{2}\,g_{N}(X_{N},Y_{N})$. we have the following:
$g(fX,fY)=g(QX,Y)-\eta(X)\eta(Y),\quad\eta(\xi)=1,\quad f(\xi)=0,\quad\eta\circ f=0,\quad\eta(X)=g(X,\xi).$
This implies that $(f,Q,\xi,\eta,g)$ is a weak almost contact metric structure on $M$. Next, we verify the weak nearly cosymplectic property on $M$ as follows:
$\nabla_{X}(fX)$ $=\nabla^{B}_{X_{B}}(F_{B}X_{B})+\nabla^{N}_{X_{N}}(J_{N}X_{N}),\quad\nabla_{X}X=\nabla^{B}_{X_{B}}X_{B}{+}\nabla^{N}_{X_{N}}X_{N}.$ (80)
Applying $f$ for both sides of the last equation gives $f\nabla_{X}X=F_{B}(\nabla_{X_{B}}X_{B})+J_{N}(\nabla_{X_{N}}X_{N}).$ Combining this with (80) and the fact that $(\nabla^{B}_{X_{B}}F_{B})X_{B}=0$ and $(\nabla^{N}_{X_{N}}J_{N})X_{N}=0$, we have
$(\nabla_{X}f)X$ $=(\nabla^{B}_{X_{B}}F_{B})X_{B}+(\nabla^{N}_{X_{N}}J_{N})X_{N}=0.$
Hence, $(f,Q,\xi,\eta,g)$ is a weak nearly cosymplectic structure on $M$. Next, since $\eta$ is pulled back from $\eta_{B}$, we have $\eta=\pi_{B}^{*}(\eta_{B})$ and $d\eta=\pi_{B}^{*}(d\eta_{B})$. Since $B^{5}=\mathbb{S}^{5}$ has a weak nearly cosymplectic structure that is not weak cosymplectic, we have $d\eta_{B}\neq 0$; hence $d\eta\neq 0$. Therefore, the structure $(f,Q,\xi,\eta,g)$ is weak nearly cosymplectic but not weak cosymplectic.
Remark 4.
For $\lambda=1$, Example 3 illustrates [14, Theorem 4.5], but for $\lambda>1$, the condition (9) is not satisfied by the weak metric structure on $M$. We do not have an example of $B^{4+s}\times\bar{M}^{\,2n-4}$ with $s>1$ illustrating Theorem 2 in the case of $s>1$. Therefore, we ask the following question: Does there exist a nontrivial weak nearly $\cal C$}-manifold ($s>1$) that is the product of a weak nearly $\cal C$}-manifold $B^{\,4+s}$ and a weak nearly Kähler manifold $N^{2n-4}$?
Conclusion
We have shown that the weak nearly $\cal C$}-structure is useful for studying metric $f$-structures, almost contact metric structures, and Killing vector fields. Some results on nearly cosymplectic manifolds (see [8, 14]) and weak nearly cosymplectic manifolds (see [18]) were generalized to weak nearly $\cal C$}-manifolds satisfying (9) and (13). Namely, the splitting Theorem 2 was proven and the $(4+s)$-dimensional case was characterized in Theorem 1. Their consequences (without the requirement (9) and (13)) present new results for nearly $\cal C$}-manifolds. Based on the numerous applications of the metric $f$-structure, we expect that the weak $\cal C$}-structure will also be useful for differential geometry (e.g., the theory of $\mathfrak{g}$-foliations, $s$-cosymplectic structures and $s$-contact structures) and theoretical physics.
Statements and Declarations
  • Funding: Sourav Nayak is financially supported by a UGC research fellowship (Grant No. 211610029330). Dhriti Sundar Patra would like to thank the Science and Engineering Research Board (SERB), India, for financial support through the Start-up Research Grant (SRG) (Grant No. SRG/2023/002264).
  • Conflict of interest/Competing interests: The authors have no conflict of interest or financial interests for this article.
  • Ethics approval: The submitted work is original and has not been submitted to more than one journal for simultaneous consideration.
  • Availability of data and materials: This manuscript has no associated data.
  • Authors’ contributions: Conceptualization, methodology, investigation, validation, writing-original draft, review, editing, and reading have been performed by all authors of the paper.