### v2

# On the splitting of weak nearly $\cal C$}-manifolds

Abstract
The interest of mathematicians in metric $f$-manifolds, in particular, almost contact metric manifolds, is motivated by the study of the geometry and dynamics of contact foliations, as well as their applications in physics. Weak metric $f$-manifolds, defined by V. Rovenski and R. Wolak (2022), open a new perspective on classical theory of $f$-manifolds and discover new applications. In this paper, we study manifolds of this type, called weak nearly $\cal C$}-manifolds, which generalize almost $\cal C$}-manifolds. We find conditions under which a $(2n+s)$-dimensional weak nearly $\cal C$}-manifold becomes locally a Riemannian product, and characterize $(4+s)$-dimensional weak nearly $\cal C$}-manifolds. The consequences of these theorems present new results for nearly $\cal C$}-manifolds.
Keywords: weak nearly $\cal C$}-manifold, weak nearly Kähler manifold, curvature tensor, totally geodesic foliation, Killing vector field.
Mathematics Subject Classifications (2010) 53C15, 53C25, 53D15
1 Introduction
Contact geometry has garnered increasing interest of mathematicians due to its significant role in physics, e.g. geometrical optics, mechanics, thermodynamics, geometric quantization, control theory, general relativity, e.g., [13, 14]. An important subclass of almost contact metric manifolds $(M^{2n+1},{f},\xi,\eta,g)$, called nearly cosymplectic manifolds, is defined in [4] by condition that the symmetric part of $\nabla{f}$ vanishes. These odd-dimensional counterparts of nearly Kähler manifolds, see [12], are used in classifying almost contact metric manifolds. Any 5-dimensional nearly cosymplectic manifold has an Einstein metric of positive scalar curvature. For example, the sphere $\mathbb{S}^{5}$ is endowed with a nearly cosymplectic structure induced by the almost Hermitian structure of $\mathbb{S}^{6}$. In dimensions greater than 5, nearly cosymplectic manifolds are locally Riemannian products ${\mathbb{R}}\times\bar{M}^{2n}$ or $B^{5}\times\bar{M}^{2n-4}$, where $\bar{M}$ is a nearly Kähler manifold and $B$ is a nearly cosymplectic manifold, see [16].
The $f$-structure introduced by K. Yano, see [24], on a smooth manifold $M^{2n+s}$ serves as a higher-dimensional analog of almost complex ($s=0$) and almost contact ($s=1$) structures. This structure is defined by a (1,1)-tensor $f$ of rank $2n$ such that $f^{3}+f=0$. The tangent bundle $TM$ splits into two complementary subbundles, $2n$-dimensional $f(TM)$ and $s$-dimensional $\ker f$. To generalize concepts and results from almost contact geometry, geometers have studied various broad classes of metric $f$-manifolds (i.e., the distribution $\ker f$ is parallelizable), e.g., [6, 7, 8]. For example, a metric $f$-manifold is termed an almost $\cal K$}-manifold if $d\Phi=0$, where $\Phi(X,Y):=g(X,{f}Y)$. The distribution $\ker f$ of a $\cal K$}-manifold is tangent to a $\mathfrak{g}$-foliation [1] with flat totally geodesic leaves. An important class of metric $f$-manifolds is given by $\cal C$}-manifolds, hence $\nabla{f}=0$, see [3]. Such manifolds are locally products of $\mathbb{R}^{s}$ and Kähler manifolds. Nearly $\cal C$}-manifolds are defined similarly to nearly Kähler manifolds, starting from $\cal C$}-manifolds, i.e., $(\nabla_{X}f)X=0$, see [2].
The question arises: Under what conditions are nearly $\cal C$-manifolds locally Riemannian products?
Recent interest in the $f$-structure among mathematicians is also motivated by the study of dynamics and integration on contact foliations, $\mathfrak{g}$-foliations, and $s$-cosymplectic manifolds, see [1, 9, 11, 15]. Contact foliations generalize to higher dimensions the flow of the Reeb vector field on contact manifolds, and $\cal K$}-structures are a particular case of uniform $s$-contact structures.
In [18], metric structures on a smooth manifold were introduced that generalize the almost Hermitian structure, the almost contact metric structure, and the metric $f$-structure. These so-called “weak metric structures” (the linear complex structure on the contact distribution is replaced by a nonsingular skew-symmetric tensor) allow us to revisit classical theory and discover new applications in the dynamics and integration on contact foliations, differential geometry of Killing vector fields, totally geodesic foliations and Ricci-type solitons, see [19], and mathematical physics, see [23]. In [20], weak nearly cosymplectic manifolds were defined, and their decomposition was proven under certain conditions, generalizing the classical result of [16]. In  [22] a special class of weak metric $f$-manifolds, called weak nearly $\cal C$-manifolds, which generalize nearly $\cal C$-manifolds, was studied.
[…]
Corollary 1.
For a nearly $\cal C$}-manifold the contact distribution is curvature invariant:
$R_{X,Y}Z\in{\cal D}\quad(X,Y,Z\in{\cal D}).$ (1)
The following consequences of Theorems 1 and 2 present new results for nearly $\cal C$}-manifolds.
Corollary 2.
Let a nearly $\cal C$-manifold $M^{\,2n+s}({f},\vec{\xi},\vec{\eta},g)$ satisfy (6) and (7). If the coframe $\vec{\eta}$ is uniform, then $n=2$ and $M$ admits an $f$-K-contact structure $(\hat{f},\vec{\xi},\vec{\eta},g)$ with $\hat{f}=-\nabla\xi_{i}$.
Corollary 3.
Let a nearly $\cal C$}- (non-$\cal C$}-) manifold $(M^{\,2n+s},{f},\vec{\xi},\vec{\eta},g)$ of $n>2$ satisfy (6) and (7). If the coframe $\vec{\eta}$ is uniform, then $M$ is locally isometric (the isometry is global if $M$ is complete simply connected) to one of the Riemannian products
[…]
the induced structure on $\bar{M}$ is nearly Kähler, and the induced structure on $B$ is a nearly $\cal C$}- structure.
2 Basic concepts of weak metric $f$-manifolds and examples
In this section, we will review some well-known concepts and results, see [18, 19, 20, 21, 22].
Recall that nearly Kähler manifolds $(M,J,g)$ are defined by A. Gray [12] using the condition that only the symmetric part of $\nabla J$ vanishes, in contrast to the Kähler case where $\nabla J=0$.
Definition 1.
A smooth manifold $M^{2n}$ of even dimension equipped with a Riemannian metric $g$ and a skew-symmetric (1,1)-tensor f{f} of rank $2n$ is called a weak almost Hermitian manifold. Such (M,f,g)(M,{f},g) is a weak Kähler manifold if $\nabla{f}=0$, where $\nabla$ is the Levi-Civita connection. A weak almost Hermitian manifold is called a weak nearly Kähler manifold, if $(\nabla_{X}{f})X=0$ for all $X\in\mathfrak{X}_{M}$.
Definition 2.
weak metric $f$-structure on a smooth manifold $M^{2n+s}$ $(n,s>0)$ is a set $({f},Q,\vec{\xi},\vec{\eta}$, $g)$, where $f$ is a skew-symmetric $(1,1)$-tensor of rank $2n$, $Q$ is a self-adjoint nonsingular $(1,1)$-tensor, $\vec{\xi}=(\xi_{1},\ldots,\xi_{s})$ are orthonormal vector fields, $\vec{\eta}=(\eta^{1},\ldots,\eta^{s})$ are dual 1-forms, and $g$ is a Riemannian metric on $M$, satisfying
${f}^{2}=-Q+\sum\nolimits_{i}{\eta^{i}}\otimes{\xi_{i}},\quad{\eta^{i}}({\xi_{j}})=\delta^{i}_{j},\quad Q\,{\xi_{i}}={\xi_{i}},$ (2)
$g({f}X,{f}Y)=g(X,Q\,Y)-\sum\nolimits_{i}{\eta^{i}}(X)\,{\eta^{i}}(Y)\quad(X,Y\in\mathfrak{X}_{M}).$ (3)
In this case, $(M^{2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ is called a weak metric $f$-manifold.
Putting $Y=\xi_{j}$ in (3), and using ${\eta^{i}}({\xi_{j}})=\delta^{i}_{j}$, we get
$\eta^{j}(X)=g(X,\xi_{j});$ (4)
thus, $\xi_{j}$} is orthogonal to the contact distribution ${\cal D}:=\bigcap_{\,i=1}^{s}\ker{\eta^{i}}$. For a more intuitive understanding of the role of $Q$ in the metric $f$-structure, we explain the following properties:
${f}\,{\xi_{i}}=0,\quad{\eta^{i}}\circ{f}=0,\quad\eta^{i}\circ Q=\eta^{i},\quad[Q,\,{f}]=0.$
By (2), $f^{2}\xi_{i}=0$ is true. From this and (2), we get $f^{3}+fQ=0$. By this, $Q\xi_{i}=\xi_{i}$ and $f^{2}\xi_{i}=0$ we get $0=-f^{3}\xi_{i}=fQ\xi_{i}=f\xi_{i}$. By $f\xi_{i}=0$, (4), and the skew-symmetry of $f$, we get $\eta^{i}(fX)=g(fX,\xi_{i})=-g(X,f\xi_{i})=0$. From this and condition ${\rm rank}\,f=2n$, we conclude that $f$ the distribution $\cal D$} of a weak metric $f$-structure is f{f}-invariant, ${\cal D}=f(TM)$ and $\dim{\cal D}=2n$. By this and $f^{3}+fQ=0$, we get $f^{3}X=f^{2}(fX)=-QfX$; hence, $f^{3}+Qf=0$. This and $f^{3}+fQ=0$ yield $fQ=Qf$. By symmetry of $Q$ and $Q\xi_{i}=\xi_{i}$, we get $\eta^{i}(QX)=g(QX,\xi_{i})=g(X,Q\xi_{i})=g(X,\xi_{i})=\eta^{i}(X)$. Therefore, the tangent bundle splits, $TM={\cal D}\oplus\ker f$ as the complementary orthogonal sum of its subbundles, the contact distribution $\cal D$}, and $\ker f=\operatorname{span}\{\xi_{1},\dots,\xi_{s}\}$ called the Reeb distribution.
The fundamental $2$-form $\Phi$ on $(M^{2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ is given by
$\Phi(X,Y)=g(X,{f}Y)\quad(X,Y\in\mathfrak{X}_{M}).$
Recall the co-boundary formulas for exterior derivative $d$ on a 1-form $\omega$} and a $2$-form $\Phi$,
$d\Phi(X,Y,Z)$ $=X\,\Phi(Y,Z)+Y\,\Phi(Z,X)+Z\,\Phi(X,Y)$
$\ -\Phi([X,Y],Z)-\Phi([Z,X],Y)-\Phi([Y,Z],X),$
$d\omega(X,Y)$ $=X({\omega}(Y))-Y({\omega}(X))-{\omega}([X,Y])\quad(X,Y\in\mathfrak{X}_{M}).$
Note that the above equality for $d\Phi$ yields
$d\Phi(X,Y,Z)=(\nabla_{X}\,\Phi)(Y,Z)+(\nabla_{Y}\,\Phi)(Z,X)+(\nabla_{Z}\,\Phi)(X,Y).$ (5)
Definition 3.
A weak metric $f$-structure is called a weak $f$-K-contact structure if $\Phi=d{\eta^{1}}=\ldots=d{\eta^{s}}$ is valid and all vector fields $\xi_{i}$} are Killing, i.e., $(\pounds_{{\xi_{i}}}\,g)(X,Y)=g(\nabla_{X}{\xi_{i}},Y)+g(\nabla_{Y}{\xi_{i}},X)=0$. Another important case of a weak metric $f$-structure is a weak almost $\cal K$}-structure, that is $d\Phi=0$. Its special case, is a weak almost $\cal C$}-structure, that is $\Phi$ and $\eta^{i}\ (1\leq i\leq s)$ are closed forms.
For $s=1$, weak almost $\cal C$}-manifolds become weak almost cosymplectic manifolds.
For weak almost $\cal K$}-manifolds (and their subclass, weak almost $\cal C$}-manifolds), the Reeb distribution $\ker f$ is tangent to a foliation. Moreover, weak almost $\cal C$}-manifolds satisfy the following conditions:
$[\xi_{i},\xi_{j}]$ $=0\quad(1\leq i,j\leq s),$ (6)
$g(\nabla_{X}\,\xi_{i},\ \xi_{j})$ $=0\quad(X\in\mathfrak{X}_{M},\ 1\leq i,j\leq s),$ (7)
trivial for $s=1$. The following condition is a consequence of (7):
$\eta^{k}(\nabla_{\xi_{i}}\,\xi_{j})=0\quad(1\leq i,j,k\leq s).$ (8)
Remark 1.
By (6), the distribution $\ker f$ of a weak almost $\cal C$}-manifold is tangent to a $\mathfrak{g}$-foliation with an abelian Lie algebra. A foliation of dimension $s$ on a smooth connected manifold $M$ is called a $\mathfrak{g}$-foliation [1], where $\mathfrak{g}$ is a Lie algebra of dimension $s$, if there exist complete vector fields $\xi_{1},\ldots,\xi_{s}$ on $M$ which, when restricted to each leaf, form a parallelism of this submanifold with a Lie algebra isomorphic to $\mathfrak{g}$.
A metric $f$-structure, satisfying (6) and $\nabla f=0$, is a normal weak almost $\cal C$}-structure, see [21, Theorem 5]. We consider the case when only the symmetric part of $\nabla f$ vanishes.
Definition 4.
A weak metric $f$-structure is called a weak nearly $\cal C$}-structure if
$(\nabla_{X}{f})Y+(\nabla_{Y}{f})X=0.$ (9)
Taking derivative of g(fV,Z)=g(V,fZ)g({f}V,Z)=-g(V,{f}Z), we see that $\nabla_{Y}{f}$ of a weak nearly $\cal C$}-manifold is skew-symmetric: $g((\nabla_{Y}{f})V,Z)=-g((\nabla_{Y}{f})Z,V)$. Taking the derivative of this, we see that $\nabla^{2}_{X,Y}{f}$ is also skew-symmetric: $g((\nabla^{2}_{X,Y}{f})V,Z)=-g((\nabla^{2}_{X,Y}{f})Z,V)$, where $\nabla^{2}_{X,Y}:=\nabla_{X}\nabla_{Y}-\nabla_{\nabla_{X}Y}$.
Example 1.
(i) Any weak Kähler manifold is obviously weak nearly Kähler. Take two or more (nearly) Kähler manifolds $(M_{j},g_{j},J_{j})$, where $J_{j}^{2}=-{\rm Id}_{\,j}$. The product $\prod_{j}(M_{j},g_{j},\sqrt{\lambda_{j}}J_{j})$, where $\lambda_{j}\neq 1$ are different positive constants, is a weak (nearly) Kähler manifold with $Q=\bigoplus_{j}\lambda_{j}{\rm Id}_{j}$. The classification of weak nearly Kähler manifolds in dimensions $\geq 4$ is an open problem.
(ii) To construct a weak nearly $\cal C$-structure $({f},Q,\vec{\xi},\vec{\eta},g)$ on the Riemannian product $M=\bar{M}\times\mathbb{R}^{s}$ of a weak nearly Kähler manifold $(\bar{M},\bar{f},\bar{g})$ with $\Omega(X,Y)=\bar{g}(X,\bar{f}Y)$ and a Euclidean space $(\mathbb{R}^{s},dy^{2})$, we take any point $(x,y)$ of $M$ and set
$\xi_{i}=(0,\partial_{\,y^{i}}),\ \ \eta^{i}=(0,dy^{i}),\ \ {f}(X,\partial_{\,y^{i}})=(\bar{f}X,0),\ \ Q(X,\partial_{\,y^{i}})=(-\bar{f}^{\,2}X,\partial_{\,y^{i}}),$
where $X\in T_{x}\bar{M}$. Note that $\nabla f=0$ if and only if $\overline{\nabla}\bar{f}=0$. On the other hand, $\overline{\nabla}\bar{f}=0$ if and only if $d\Omega=0$, see (5) with $\Phi=\Omega$, i.e., $(\bar{M},\Omega)$ is a symplectic manifold.
(iii) Let a Riemannian manifold $(M^{2n+s},g)$ admit two nearly $\mathcal{C}$}-structures with common vector fields $\xi_{i}$ and one-forms $\eta^{i}=g(\xi_{i},\cdot)$ for $1\leq i\leq s$. Let $f_{1}\neq f_{2}$ be such that the self-adjoint tensor $\psi:=f_{1}f_{2}+f_{2}f_{1}$ is not identically zero. Then the tensor $f:=(\cos t)\,f_{1}+(\sin t)\,f_{2}$ for small constant $t>0$ satisfies the condition (9). Observe that $f^{2}=-Q+\sum\nolimits_{i}\eta^{i}\otimes\xi_{i}$ with $Q={\rm Id}-(\sin t\cos t)\,\psi$, which is positive definite for small $t>0$. Thus, $(f,Q,\vec{\xi},\vec{\eta},g)$ is a weak nearly $\cal C$-structure on $M^{2n+s}$.
Let denote the $\cal D$} component of a vector. The following condition is trivial when $Q={\rm Id}$:
$((\nabla_{X}Q)Y)^{\top}=0,\quad(\nabla_{Y}Q)Y=0\quad(X\in TM,\ Y\in{\cal D}).$ (10)
Using the first equality of (10), for any $X,Y\in TM$, we have
$(\nabla_{X}Q)Y$ $=[(\nabla_{X}Q)Y]^{\top}+\sum\nolimits_{i}\eta^{i}((\nabla_{X}Q)Y)\,\xi_{i}$
$=-\sum\nolimits_{i}\{\eta^{i}(Y)\widetilde{Q}\nabla_{X}\xi_{i}+g(\widetilde{Q}\nabla_{X}\xi_{i},Y)\,\xi_{i}\}.$ (11)
Remark 2.
In [19], instead of the overly strict condition (4), the condition (10) should be used.
Lemma 1.
On a weak nearly $\cal C$}-manifold satisfying (6) and (8), the Reeb distribution $\ker f$ defines a flat totally geodesic foliation; moreover, if (7) and (10) hold, then the vector fields $\xi_{i}$ are Killing.
Proof.
The first part easily follows from [22, Proposition 2]. For the second part, following the proof of [22, Proposition 2], we have $(\mathcal{L}_{\xi_{j}}\,g)(\xi_{k},\cdot)=0$ and
$(\mathcal{L}_{\xi_{j}}\,g)(X,Y)=\eta^{j}\big((\nabla_{X}Q)Y+(\nabla_{Y}Q)X\big)\qquad(X,Y\in{\cal D}).$
By this and the second equality of (10), we obtain $(\mathcal{L}_{\xi_{j}}\,g)(X,Y)=0$ for $X,Y\in{\cal D}$. Combining this with $(\mathcal{L}_{\xi_{j}}\,g)(\xi_{k},\cdot)=0$, we conclude that $\xi_{j}$ is Killing. ∎
For a Riemannian manifold $(M,g)$ equipped with a Killing vector field $\zeta$}, we get, see [24],
$\nabla_{X}\nabla_{Y}\,{\zeta}-\nabla_{\nabla_{X}Y}\,{\zeta}=R_{\,X,\,{\zeta}}\,Y,$ (12)
where $R_{{X},{Y}}=[\nabla_{X},\nabla_{Y}]-\nabla_{[X,Y]}$ is the curvature tensor. We have $g(R_{\,\xi,Z}\,{f}X,{f}Y)=0$ for a nearly cosymplectic manifold, see [10]; thus its distribution ${\cal D}=\ker\eta$ is curvature invariant, see (1). In our study of weak nearly $\cal C$}-manifolds, we assume a condition weaker than (1):
$R_{\widetilde{Q}X,Y}Z\in{\cal D}\quad(X,Y,Z\in{\cal D}).$ (13)
Obviously, conditions (10) and (13) become trivial for metric $f$-manifolds, i.e., when $Q={\rm Id}$.
3 Main results
Here, we prove a splitting result (Theorem 1) for weak nearly $\cal C$}-manifolds satisfying some additional conditions such as (10) and (13). Theorem 2 characterizes $(4+s)$-dimensional weak nearly $\cal C$}-manifolds.
We define the tensor field ${\bf h}=(h_{1},\ldots,h_{s})$, as in the classical case, e.g., [10], where
$h_{i}=\nabla\xi_{i}\quad(1\leq i\leq s).$ (14)
As a consequence of Lemma 1, the tensor fields $h_{i}$ are skew-symmetric:
$g(h_{i}X,\,X)=g(\nabla_{X}\,\xi_{i},X)=\tfrac{1}{2}\,(\pounds_{\xi_{i}}\,g)(X,X)=0\quad(1\leq i\leq s).$
Using (7), we get $\eta^{j}\circ h_{i}=0$ and $h_{i}\,\xi_{j}=0$ for $1\leq i,j\leq s$. This shows that $\operatorname{dim(ker}h_{i})\geq s$ for all $i$.
The following result generalizes [20, Proposition 4.1] (and [16, Proposition 4.2]) to the case $s>1$.
Proposition 1.
For a weak nearly $\cal C$}-manifold with conditions (6), (7), (10) and (13), the operators $h_{i}$ and $h_{j}$ commute and each composition $h_{i}h_{j}$ is self-adjoint for $1\leq i,j\leq s$. Moreover, the eigenvalues and their multiplicities of $h_{i}h_{j}$ and $h_{i}^{2}$ are constant, and $h_{i}h_{j}$ and $h_{i}^{2}$ share the same eigen-frame.
Proof.
From the first Bianchi identity, using (31), we have $R_{\xi_{i},X}\,\xi_{j}=R_{\xi_{j},X}\,\xi_{i}$ for all $X\in\mathfrak{X}_{M}$ and $1\leq i,j\leq s$. From this, applying (51) and (50), we get the commutativity of $h_{i}$ and $h_{j}$:
$0=R_{\xi_{i},X}\,\xi_{j}-R_{\xi_{j},X}\,\xi_{i}=-(\nabla_{X}h_{i})\xi_{j}+(\nabla_{X}h_{j})\xi_{i}=h_{i}h_{j}X-h_{j}h_{i}X.$
The first part follows from the commutativity and skew-symmetry of the operators $h_{i}\ (1\leq i\leq s)$.
From the above and Lemma 5 we obtain
$(\nabla_{X}\,h_{i}h_{j})Y=h_{i}(\nabla_{X}\,h_{j})Y+(\nabla_{X}\,h_{i})h_{j}Y=\sum\nolimits_{k}\big[g(X,h_{k}h_{i}h_{j}Y)\,\xi_{k}-\eta^{j}(Y)\,h_{i}h_{j}h_{k}X\big].$ (15)
Consider an eigenvalue $\mu$ of $h_{i}h_{j}$ and a local unit vector field $Y\in\mathcal{D}$ such that $h_{i}h_{j}Y=\mu Y$. Applying (15) for any nonzero vector fields $X,Y\in\mathcal{D}$, we find $g((\nabla_{X}\,h_{i}h_{j})Y,Y)=0$, thus
$0=g((\nabla_{X}\,h_{i}h_{j})Y,Y)=g(\nabla_{X}\,(h_{i}h_{j}Y),Y)-g(h_{i}h_{j}(\nabla_{X}\,Y),Y)$
$\ \ =X(\mu)\,g(Y,Y)+\mu\,g(\nabla_{X}\,Y,Y)-g(\nabla_{X}\,Y,h_{i}h_{j}Y)=X(\mu)\,g(Y,Y),$
which implies that $X(\mu)=0$ for all $X\in\mathfrak{X}_{M}$. When $j=i$, we get a similar result for $h_{i}^{2}$.
Next, by the above, the operators $h_{i}$ and $h_{j}$ commute for all $i,j$, which implies that $h_{i}^{2}$ and $h_{j}^{2}$ also commute. Each $h_{i}^{2}$ is self-adjoint, and therefore, diagonalizable.
It is well known that a finite family of diagonalizable operators on a finite-dimensional Euclidean space is simultaneously diagonalizable if and only if the opeators commute. By the above, there exists an orthonormal basis $\{e_{k}\}$ of $TM$ such that $h_{i}^{2}e_{k}=\lambda_{i,k}\,e_{k}$ for all $1\leq i\leq s$ and $k$. ∎
By Proposition 1, the spectrum of the self-adjoint operator $h_{i}^{2}$ has the form
$Spec(h_{i}^{2})=\{0,-\lambda_{i,1}^{2},\ldots-\lambda_{i,r_{i}}^{2}\}\qquad(1\leq i\leq s),$ (16)
where $r_{i}\geq 1$, $\lambda_{i,j}$ is a positive real number and $\lambda_{i,j}\neq\lambda_{i,k}$ for $j\neq k$. If $X\neq 0$ is an eigenvector of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$, then $X,{f}X,h_{i}X$ and $h_{i}\,{f}X$ are orthogonal nonzero eigenvectors of $h_{i}^{2}$ with the same eigenvalue $\lambda^{2}_{i,j}$. Since $h_{i}(\xi_{j})=0$ for all $1\leq i,j\leq s$, the Reeb distribution $\ker f=\operatorname{span}\{\xi_{1},\dots,\xi_{s}\}$ is contained in the kernel of $h_{i}$. Consequently, the eigenvalue 0 of $h_{i}^{2}$ has multiplicity $2p_{i}+s$ for some integer $p_{i}\geq 0$, associated with each operator $h_{i}^{2}$. Denote by
$\bullet$ $D_{i,0}$ the smooth distribution of the eigenvectors of $h_{i}^{2}$ with eigenvalue 0 orthogonal to $\ker f$;
$\bullet$ $D_{i,j}\ (1\leq j\leq r_{i})$ be the smooth distribution of the eigenvectors of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$.
Note that $D_{i,0}$ and $D_{i,j}$ belong to $\cal D$} and are f{f}- and $h_{k}$-invariant for each $1\leq i,k\leq s$.
Remark 3.
By the skew-symmetry of $h_{i}$, for every $X,Y\in\mathfrak{X}_{M}$, we have
$d\eta^{i}(X,Y)=2\,g(h_{i}X,Y)\quad(1\leq i\leq s).$ (17)
The above shows that $\operatorname{ker}h_{i}=\operatorname{ker}d\eta^{i}$ and as a consequence of this, $\operatorname{dim\,(ker}d\eta^{i})\geq s$ for all $i$.
The following result generalizes [20, Proposition 4.3] (and [16, Proposition 4.1]) to the case $s>1$.
Proposition 2.
[…]
Proof.
By (17) and the condition ${\bf h}\equiv 0$, we have $d\eta^{i}=0$ for all $i$. We have
$0=d\eta^{i}(X,Y)=-\eta^{i}([X,Y])\quad(1\leq i\leq s,\ X,Y\in{\cal D}),$
which indicates that $X,Y]\in{\cal D}$; thus the distribution $\cal D$} is integrable. Integral manifolds of $\cal D$} are totally geodesic as
$g(\nabla_{X}\,Y,\xi_{i})=-g(Y,h_{i}X)=0\quad(1\leq i\leq s,\ X,Y\in{\cal D}).$
Since $\operatorname{ker}f$ is integrable, totally geodesic and flat (see Lemma 1), by de Rham Decomposition Theorem, the manifold is locally the Riemannian product $\mathbb{R}^{s}\times\bar{M}$. The weak nearly $\cal C$}-structure on $M$ induces on $\bar{M}$ a weak nearly Kähler structure.
Conversely, if a weak nearly $\cal C$}-manifold is locally the Riemannian product $\mathbb{R}^{s}\times\bar{M}$, where $\bar{M}$ is a weak nearly Kähler manifold and $\xi_{i}=(\partial_{x_{i}},0)$ ($(x,y)\in\mathbb{R}^{s}\times\bar{M}$), then $d\eta^{i}=0$ for $1\leq i\leq s$ or simply ${\bf h}\equiv 0$, by (17). By this and the definition of $\xi_{i}$, the conditions (6) and (7) are true. ∎
The following result generalizes [20, Proposition 4.2] (and [16, Proposition 4.3]) to the case $s>1$.
Proposition 3.
Let $(M^{\,2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ be a weak nearly $\mathcal{C}$-manifold with conditions (6), (7), (10) and (13). Then
  • (i)
    distributions $\ker f\oplus{\cal D}_{i,j}\ (1\leq i\leq s,\,1\leq j\leq r_{i})$ are integrable with totally geodesic leaves.
Moreover, if $\dim(\ker d\eta^{i})>s$ for all $i$, then the following is true:
  • (ii)
    ${\cal D}_{i,0}\neq 0$ for all $i$, and distributions $\operatorname{ker}f\oplus{\cal D}_{i,0}$ are integrable with totally geodesic leaves;
  • (iii)
    if the exterior derivatives of the coframe $\vec{\eta}$ have a common kernel, then distributions ${\cal D}_{i,0}$ coincide, are integrable with totally geodesic leaves, and the induced structure on each leaf of ${\cal D}_{i,0}$ is a weak nearly Kähler structure $(\bar{f},\bar{g})$ with the property $\bar{\nabla}(\bar{f}^{\,2})=0$;
  • (iv)
    the distribution $\ker f\oplus{\cal D}_{i,1}\oplus\ldots\oplus{\cal D}_{i,r_{i}}$ is integrable with totally geodesic leaves.
Proof.
Consider an eigenvector $X$ of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$, see (16). Then $\nabla_{X}\,\xi_{k}=h_{k}X\in{\cal D}_{i,j}$ ($k=1,\ldots,s$), as $h_{i}^{2}(h_{k}X)=h_{k}(h_{i}^{2}X)=-\lambda^{2}_{i,j}\,h_{k}X$. Since (15) implies $\nabla_{\xi_{k}}h_{i}^{2}=0$, then $\nabla_{\xi_{k}}X$ is also an eigenvector of $h_{i}^{2}$ with eigenvalue $\lambda^{2}_{i,j}$ for all $k=1,..,s$. By Lemma 1, the distribution $\operatorname{ker}f$ defines a flat totally geodesic foliation. Moreover, since $\nabla_{X}\,\xi_{k},\ \nabla_{\xi_{k}}X\in\mathcal{D}_{i,j}$ for all $X\in\mathcal{D}_{i,j}$, we have $\operatorname{ker}f\oplus\mathcal{D}_{i,j}$ is integrable (involutive). Now, taking $X,Y\in{\cal D}_{i,j}$ and applying (15), we get
$h_{i}^{2}(\nabla_{X}\,Y)=-\lambda^{2}_{i,j}\,\nabla_{X}Y-(\nabla_{X}\,h_{i}^{2})Y=-\lambda^{2}_{i,j}\,\nabla_{X}Y+\lambda^{2}_{i,j}\sum\nolimits_{k}g(X,h_{i}Y)\,\xi_{k}.$
Using the above and the anti-commutativity of $f$ and $h_{i}$, we have
$h_{i}^{2}({f}^{2}\nabla_{X}Y)={f}^{2}(h_{i}^{2}\nabla_{X}Y)=-\lambda^{2}_{i,j}\,{f}^{2}(\nabla_{X}Y);$
hence, ${f}^{2}\nabla_{X}Y\in{\cal D}_{i,j}$. Similarly, using (26), we get $\widetilde{Q}\nabla_{X}Y\in{\cal D}_{i,j}$. It follows that
$\nabla_{X}Y=-\widetilde{Q}\,\nabla_{X}Y-{f}^{2}\nabla_{X}Y+\sum\nolimits_{k}\eta^{k}(\nabla_{X}Y)\,\xi_{k},$
see (2), belongs to the distribution $\ker f\oplus{\cal D}_{i,j}$. This proves $(i)$.
[…]
Since $\mathcal{D}_{i,0}\neq 0$, we get $h_{i}^{\,2}(\nabla_{X}Y)=0$ for all $X,Y\in\mathcal{D}_{i,0}$. In addition, we compute
$g(\nabla_{X}Y,\xi_{k})=-g\big(Y,\nabla_{X}\xi_{k}\big)=-g(Y,h_{k}X),$
which vanishes if $\operatorname{ker}h_{i}=\operatorname{ker}h_{k}$ for all $1\leq i,k\leq s$. By (17), this condition is equivalent to the fact that the exterior derivatives of the coframe $\vec{\eta}$ have a common kernel. Thus, each distribution $D_{i,0}$ defines a totally geodesic foliation if and only if the exterior derivatives of the coframe $\vec{\eta}$ have a common kernel. By (9) and Lemma 2, the leaves of $D_{i,0}$ are f{f}-invariant and $Q$-invariant. Thus, the weak nearly $\cal C$}-structure on $M$ with conditions (6), (7), (10) and (13) induces a weak nearly Kähler structure $(\bar{f},\bar{g})$ on each leaf of $D_{i,0}$ with the property $\bar{\nabla}(\bar{f}^{\,2})=0$, where $\bar{\nabla}$ is the Levi-Civita connection of $\bar{g}$. This proves $(iii)$.
To establish $(iv)$, it suffices to verify that for $X\in{\cal D}_{i,j}$ and $Y\in{\cal D}_{i,k}$ with $j\neq k$, one has $\nabla_{X}Y\in\ker f\oplus{\cal D}_{i,1}\oplus\cdots\oplus{\cal D}_{i,r}$. Similarly to part $(i)$, we obtain
$h_{i}^{2}(\nabla_{X}Y)=-\lambda_{i,k}\,\nabla_{X}Y+\lambda_{i,k}\sum\nolimits_{j}g(X,h_{j}Y)\,\xi_{j};$
hence, $h_{i}^{2}(f^{2}\nabla_{X}Y)=-\lambda_{i,k}\,f^{2}\nabla_{X}Y$. Thus $f^{2}\nabla_{X}Y\in{\cal D}_{i,k}$, and we calculate $\widetilde{Q}\nabla_{X}Y\in{\cal D}_{i,k}$. Combining these observations, we may write
$\nabla_{X}Y=-\widetilde{Q}(\nabla_{X}Y)-f^{2}\nabla_{X}Y+\sum\nolimits_{k=1}^{s}\eta^{k}(\nabla_{X}Y)\,\xi_{k}\in\operatorname{ker}f\oplus{\cal D}_{i,k}\subset\ker f\oplus{\cal D}_{i,1}\oplus\cdots\oplus{\cal D}_{i,r}.$
Similarly, we obtain $\nabla_{Y}X\in\operatorname{ker}f\oplus{\cal D}_{i,j}$. Consequently, both $[X,Y]$ and $\nabla_{X}Y$ belong to $\ker f\oplus{\cal D}_{i,1}\oplus\cdots\oplus{\cal D}_{i,r}$, which completes the proof. ∎
Definition 5 (see [11]).
Let $M$ be a $(2n+s)$-dimensional smooth manifold where $n,s$ are positive integers. An $s$-contact structure on $M$ is a collection $\vec{\eta}=(\eta^{1},\cdots,\eta^{s})$ of $s>1$ pointwise linearly independent non-vanishing one-forms, together with a splitting $TM={\mathcal{R}}\oplus{\cal D}$ of the tangent bundle, satisfying the following conditions:
  • (i)
    ${\cal D}:=\bigcap_{\,i}\ker\eta^{i}$;
  • (ii)
    $d\eta^{i}|_{\cal D}$ is non-degenerate for every $i$;
  • (iii)
    $\ker d\eta^{i}={\mathcal{R}}$ for every $i$.
A manifold $M$ endowed with such a structure is called an $s$-contact manifold and is denoted by $(M,\vec{\eta},{\mathcal{R}}\oplus{\cal D})$. A coframe $\vec{\eta}$ (and the $s$-contact structure it defines) is said to be uniform if it satisfies
  • (iv)
    $d\eta^{i}=d\eta^{j}$ for all $1\leq i,j\leq s$.
An example of $s$-contact manifolds is a weak $f$-K-contact manifold, in fact, it is uniform. By (17), $d\eta^{i}=d\eta^{j}\iff h_{i}=h_{j}$ for all $1\leq i,j\leq s$.
Next, we will use Lemmas 37 to characterize $(4+s)$-dimensional weak nearly $\cal C$}-manifolds.
The following result generalizes [20, Theorem 4.2] (see also Theorem 4.4 in [16]) to the case $s>1$.
Theorem 1.
Let $(M^{\,2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ be a weak nearly $\cal C$-manifold satisfying (6), (7), (10) and (13). If the coframe $\vec{\eta}$ has a common exterior derivative, see (iv) of Definition 5, then $n=2$ and $M$ admits a weak $f$-K-contact structure $(\hat{f},\hat{Q},\vec{\xi},\vec{\eta},g)$ defined by $\hat{f}=-\nabla\xi_{i}$ and $\hat{Q}X=R_{\xi_{i}}X\ (X\in{\cal D})$.
Proof.
We consider 2-forms $\Phi^{(k)}_{i}(X,Y)=g({f}h_{i}^{k}X,Y)$ and $\Phi^{(1)}_{i,j}(X,Y)=g({f}h_{i}h_{j}X,Y)$, where $i=1,\dots s,\ k=0,1$. Also, since $h_{i}$ and $h_{j}$ commute, we have $\Phi^{(1)}_{i,j}=\Phi^{(1)}_{j,i}$. It is easy to calculate $d\Phi(X,Y,Z)=g((\nabla_{X}{f})Z,Y)+g((\nabla_{Y}\,{f})X,Z)+g((\nabla_{Z}\,{f})Y,X)$, see [17]. We will show that
$d\Phi^{(0)}_{i}=3\sum\nolimits_{j}\eta^{j}\wedge\Phi^{(1)}_{j},\quad d\Phi^{(1)}_{i}=3\sum\nolimits_{j}\eta^{j}\wedge\Phi^{(1)}_{i,j}$. (18)
Indeed, applying (14) and ${f}\,\xi_{i}=0$, we find the $\xi_{i}$-component of $(\nabla_{X}{f})Y$:
$g((\nabla_{X}{f})Y,\xi_{i})=-g((\nabla_{X}{f})\,\xi_{i},Y)=g({f}\nabla_{X}\,\xi_{i},Y)=g({f}h_{i}X,Y).$ (19)
Replacing $Z$ by fZ{f}Z in (57) and using (9), we obtain for all $1\leq i\leq s$
$-g((\nabla_{X}{f})Y,{f}^{2}Qh_{j}Z)=\sum\nolimits_{i}\big[$ $\eta^{i}(X)g(h_{j}Y,h_{i}fZ){-}\eta^{i}(Y)g(h_{j}X,h_{i}fZ)$
$+\eta^{i}(X)g(\widetilde{Q}(Q{+}{\rm Id})Y,h_{i}h_{j}fZ)\big].$ (20)
By conditions, ${\bf h}\not\equiv 0$ on ${\cal D}\setminus\{0\}$ and $Q(\mathcal{D})=\mathcal{D}$, thus from (3) we get $g((\nabla_{X}\,{f})Y,V)=0$ for $X,Y,V\in{\cal D}$. By the above and (19), using $X=X^{\top}+\sum\nolimits_{i}\eta^{i}(X)\,\xi_{i}$ and $Y=Y^{\top}+\sum\nolimits_{i}\eta^{i}(Y)\,\xi_{i}$, we obtain
$g((\nabla_{X}\,{f})Y,V)$
$=\sum\nolimits_{i}\big[\eta^{i}(V)\,g((\nabla_{X^{\top}}\,{f})Y^{\top},\xi_{i})+\eta^{i}(X)\,g((\nabla_{\,\xi_{i}}\,{f})Y^{\top},V)+\eta^{i}(Y)\,g((\nabla_{X^{\top}}\,{f})\,\xi_{i},V)\big]$
$=\sum\nolimits_{i}\big[-\eta^{i}(V)\,g((\nabla_{X^{\top}}\,{f})\,\xi_{i},Y^{\top})-\eta^{i}(X)\,g((\nabla_{Y^{\top}}\,{f})\,\xi_{i},V)+\eta^{i}(Y)\,g((\nabla_{X^{\top}}\,{f})\,\xi_{i},V)\big]$
$=\sum\nolimits_{i}\big[\eta^{i}(V)\,g({f}h_{i}X,Y)+\eta^{i}(X)\,g({f}h_{i}Y,V)+\eta^{i}(Y)\,g({f}h_{i}V,X)\big].$ (21)
Using the expression for $d\Phi_{i}^{(0)}$ and (3), we have the first equality of (18).
Similarly, using (3) and (53) of Lemma 5, we get
$g((\nabla_{X}\,({f}\,h_{i}))Y,Z)=g((\nabla_{X}{f})h_{i}Y,Z)+g({f}(\nabla_{X}\,h_{i})Y,Z)$
$=\sum\nolimits_{j}\big[\eta^{j}(X)\,g({f}\,h_{i}h_{j}Y,Z)+\eta^{j}(Y)\,g({f}\,h_{i}h_{j}Z,X)+\eta^{j}(Z)\,g({f}\,h_{i}h_{j}X,Y)\big],$
which implies the second equality of (18). From (18), using the symmetry: $\Phi^{(1)}_{i,j}=\Phi^{(1)}_{j,i}$ and the anti-symmetry of wedge products: $\eta^{j}\wedge\eta^{i}=-\eta^{i}\wedge\eta^{j}$, we obtain
$0=d^{2}\Phi^{(0)}_{i}=3\sum\nolimits_{j}\big[d\eta^{j}\wedge\Phi^{(1)}_{j}-\eta^{j}\wedge d\Phi^{(1)}_{j}\big]=3\sum\nolimits_{j}d\eta^{j}\wedge\Phi^{(1)}_{j}$. (22)
Substituting $Y=\xi_{i}$ into (53) and applying (23), we represent the $\xi$-sectional curvature as
$K(\xi_{i},X)=g(h_{i}X,h_{i}X)=\|h_{i}X\|^{2}\quad(X\in\mathcal{D},\ g(X,X)=1,\ 1\leq i\leq s).$
Under the assumptions, the $\xi$-sectional curvature of $(M,g)$ is positive, which implies that the Jacobi operators $R_{\,{\xi_{i}}}:X\mapsto R_{X,\,{{\xi_{i}}}}\,{{\xi_{i}}}\ (1\leq i\leq s)$ are positive definite on $\mathcal{D}$. By [19, Theorem 2], there exists a weak $f$-K-contact structure $(\hat{f},\hat{Q},\vec{\xi},\vec{\eta},g)$ on $M$ defined by $\hat{f}=-\nabla\xi_{i}$ and $\hat{Q}X=R_{\xi_{i}}X\ (X\in{\cal D})$.
As a consequence of this, $\eta^{i}$ is a contact form and $d\eta^{i}$ is non-degenerate on $\mathcal{D}$. Thus, $M$ is a uniform $s$-contact manifold, see Definition 5. Now, if $\dim M>4+s$, then (since $M$ is a uniform $s$-contact manifold) (22) implies that $d\eta^{j}\wedge\sum_{j}\Phi^{(1)}_{j}=0\,\text{for all }j.$ By Lemma 7, this condition forces $\Phi^{(1)}_{i}=0$ for all $i$; hence, ${\bf h}\equiv 0$. This contradicts the fact that $d\eta^{i}$ is non-degenerate on $\mathcal{D}$. Therefore, $\dim M\leq 4+s$. Finally, since $d\eta^{i}$ is non-degenerate on $\mathcal{D}$, the distribution $\operatorname{ker}d\eta^{i}$ has dimension $s$. We conclude that $h_{i}^{2}|_{\cal D}$ is non-zero and has at least one non-zero eigenvalue, say, $\lambda^{2}$, whose multiplicity is $4$: if $X$ is an eigenvector of $h_{i}^{2}$ with eigenvalue $\lambda^{2}$, then $\{X,fX,h_{i}X,h_{i}fX\}$ are also orthogonal eigenvectors of $h_{i}^{2}$ with the same eigenvalue. Therefore, $M$ is a $(4+s)$-dimensional manifold and the multiplicity of the eigenvalue $\lambda^{2}$ of $h_{i}^{2}$ is $4$. ∎
Remark 4.
The weak $f$-K-contact structure in Theorem 1 does not coincide with the given weak nearly $\mathcal{C}$-structure $({f},Q,\vec{\xi},\vec{\eta},g)$, even though the $1$-forms $\eta^{i}$ are identical. Nevertheless, each $\eta^{i}$ is a contact form, thus $d\eta^{i}$ is non-degenerate on $\mathcal{D}$.
We generalize the splitting Theorem 4.1 in [20] (see also Theorem 4.5 in [16]) to the case $s>1$.
Theorem 2.
[…]
$(i)~\mathbb{R}^{s}\times\bar{M}^{\,2n},\quad(ii)~B^{\,4+s}\times\bar{M}^{\,2n-4},$
[…]
Proof.
If ${\bf h}\equiv 0$, then by Proposition 2, $M$ is locally isometric to $\mathbb{R}^{s}\times\bar{M}^{\,2n}$.
[…]
Remark 5.
In aim to obtain an example illustrating Theorem 2, we ask the following question: Does there exist a nontrivial weak nearly $\cal C$}-manifold $B^{\,4+s}$ with $s>1$ ?
4 Auxiliary results
In this section, we consider weak nearly $\mathcal{C}$-manifolds satisfying some of conditions (6), (7), (10) and (13). We prove lemmas generalizing known properties of weak nearly cosymplectic manifolds.
The following result generalizes [20, Lemma 3.1] for the case of $s>1$.
Lemma 2.
For a weak nearly $\cal C$}-manifold $(M^{\,2n+s},{f},Q,\vec{\xi},\vec{\eta},g)$ satisfying (6) and (7), we get
$(\nabla_{X}\,h_{i})\,\xi_{j}=-h_{i}h_{j}X\quad(1\leq i,j\leq s),$ (23)
$(\nabla_{X}{f})\,\xi_{i}=-{f}\,h_{i}X\quad(1\leq i\leq s),$ (24)
$h_{i}\,{f}+{f}\,h_{i}=0\quad(1\leq i\leq s)\quad(h_{i}\ {\rm anticommutes\ with}\ {f}),$ (25)
$h_{i}\,Q=Q\,h_{i}\quad(1\leq i\leq s)\quad(h_{i}\ {\rm commutes\ with}\ Q).$ (26)
Proof.
Differentiating the equality $h_{i}\,\xi_{j}=0$ and using (14), we obtain (23):
$0=\nabla_{X}\,(h_{i}\,\xi_{j})=(\nabla_{X}\,h_{i})\,\xi_{j}+h_{i}(\nabla_{X}\,\xi_{j})=(\nabla_{X}\,h_{i})\,\xi_{j}+h_{i}h_{j}X.$
Equation (24) is obtained by differentiating the relation $f\xi_{i}=0$ and using the definition of $h_{i}$. Differentiating $g({f}\,Y,\xi_{i})=0$ yields $0=g((\nabla_{X}{f})Y,\xi_{i})+g({f}\,Y,h_{i}X)$. Interchanging $X$ and $Y$ in the last equation, then adding the resulting expression to the original and applying (9), yields (25). Finally, from (2) and (7) we get $f^{2}h_{i}X-h_{i}f^{2}X=-Qh_{i}X+h_{i}QX$ for all $X\in\mathfrak{X}_{M}$, and (26) follows from this by applying (25). ∎
The following lemmas generalize certain formulas in [20, Lemmas 3.2-3.5] for the case of $s>1$.
Lemma 3.
For a weak nearly $\cal C$}-manifold satisfying (10) we obtain
$g((\nabla_{X}{f}){f}Y,Z)$ $=g((\nabla_{X}{f})Y,{f}Z)+\sum\nolimits_{i}\big[\eta^{i}(Y)g(h_{i}X,QZ)+\eta^{i}(Z)g(h_{i}X,QY)\big],$ (27)
$g((\nabla_{{f}X}{f})Y,Z)$ $=g((\nabla_{X}{f})Y,{f}Z)+\sum\nolimits_{i}\big[\eta^{i}(X)g(h_{i}Z,QY)+\eta^{i}(Z)g(h_{i}X,QY)\big],$ (28)
$g((\nabla_{{f}X}{f}){f}Y,Z)$ $=g((\nabla_{X}{f})QZ,Y)+\sum\nolimits_{i}\big[\eta^{i}(X)g(h_{i}Z,QfY)+\eta^{i}(Y)g(h_{i}X,QfZ)$
$\ +\eta^{i}(Z)g({f}h_{i}X,\widetilde{Q}Y)\big].$ (29)
Proof.
Differentiating (2) and using (2), and the skew-symmetry of $\nabla_{X}{f}$, we get (27). We obtain (28) from (27) by the definition of weak nearly $\cal C$}-manifolds, followed by switching $X$ and $Y$. Replacing $Y$ by fY{f}Y in (28) and using (27) and (2), we get (3). ∎
Lemma 4.
The curvature tensor of a weak nearly $\cal C$}-manifold satisfies the equalities
$g(R_{{f}X,Y}Z,V)+g(R_{X,{f}Y}Z,V)+g(R_{X,Y}{f}Z,V)+g(R_{X,Y}Z,{f}V)=0,$ (30)
$R_{\,\xi_{j},\,\xi_{k}}=0\quad(1\leq j,k\leq s);$ (31)
moreover, if the condition (13) is true, then
$g(R_{\,\xi_{j},Z}\,{f}X,{f}Y)=0\quad(1\leq j\leq s),\quad\mbox{\rm hence,}\ \mathcal{D}\ \mbox{\rm is curvature invariant}.$ (32)
Proof.
Differentiating (9), we find
$(\nabla^{2}_{X,Y}\,{f})Z+(\nabla^{2}_{X,Z}\,{f})Y=0\quad(X,Y,Z\in\mathfrak{X}_{M}).$ (33)
Applying the Ricci identity, see [10],
$g((\nabla^{2}_{X,Y}{f})V,Z)-g((\nabla^{2}_{Y,X}{f})V,Z)=g(R_{X,Y}{f}V,Z)+g(R_{X,Y}V,{f}Z),$ (34)
from (33) and the skew-symmetry of $\nabla^{2}_{X,Y}\,{f}$ we get
$g(R_{X,Y}Z,{f}V)-g(R_{X,Y}V,{f}Z)+g((\nabla^{2}_{X,Z}\,{f})Y,V)-g((\nabla^{2}_{Y,Z}\,{f})X,V)=0.$ (35)
By Bianchi and Ricci identities, we find
$g(R_{X,Y}Z,{f}V)$ $=-g(R_{Y,Z}X,{f}V)-g(R_{Z,X}Y,{f}V)$
$=g((\nabla^{2}_{Y,Z}\,{f})V,X)-g((\nabla^{2}_{Z,Y}\,{f})V,X)-g(R_{Y,Z}V,{f}X)-g(R_{Z,X}Y,{f}V).$ (36)
Substituting (4) into (35), we get
$g(R_{X,Z}Y,{f}V)-g(R_{X,Y}V,{f}Z)-g(R_{Y,Z}V,{f}X)$
$-g((\nabla^{2}_{Z,Y}\,{f})V,X)-g((\nabla^{2}_{X,Z}\,{f})V,Y)=2\,g((\nabla^{2}_{Y,Z}\,{f})X,V).$ (37)
On the other hand, using the Ricci identity (34) again, we see that
$g(R_{X,Z}Y,{f}V)-g(R_{X,Z}V,{f}Y)-g((\nabla^{2}_{X,Z}\,{f})Y,V)+g((\nabla^{2}_{Z,X}\,{f})Y,V)=0.$ (38)
Adding (38) to (4) and applying (33), we get
$2\,g(R_{X,Z}Y,{f}V)-g(R_{X,Y}V,{f}Z)-g(R_{Y,Z}V,{f}X)-g(R_{X,Z}V,{f}Y)=2\,g((\nabla^{2}_{Y,V}\,{f})Z,X).$ (39)
Swapping $Y$ and $V$ in (39), we find
$2\,g(R_{X,Z}V,{f}Y)-g(R_{X,V}Y,{f}Z)-g(R_{V,Z}Y,{f}X)-g(R_{X,Z}Y,{f}V)=2\,g((\nabla^{2}_{V,Y}\,{f})Z,X).$ (40)
Subtracting (40) from (39), and using Bianchi and Ricci identities, we get the equality, which by replacing $Z$ and $Y$ gives (30).
Replacing $X$ by fX{f}X in (30) and using (2), we have
$-g(R_{QX,Y}Z,V)$ $+\sum\nolimits_{i}\eta^{i}(X)\,g(R_{\,\xi_{i},Y}Z,V)+g(R_{{f}X,{f}Y}Z,V)$
$+g(R_{{f}X,Y}{f}Z,V)+g(R_{{f}X,Y}Z,{f}V)=0.$ (41)
Exchanging $X$ and $Y$ in (4), we find
$g(R_{X,QY}Z,V)$ $+\sum\nolimits_{i}\eta^{i}(Y)\,g(R_{\,\xi_{i},X}Z,V)-g(R_{{f}X,{f}Y}Z,V)$
$+g(R_{{f}Y,X}{f}Z,V)+g(R_{{f}Y,X}Z,{f}V)=0.$ (42)
Subtracting (4) from (4), we obtain
$2\,g(R_{{f}X,{f}Y}Z,V)-2\,g(R_{X,Y}Z,V)+\sum\nolimits_{i}[\eta^{i}(X)\,g(R_{\,\xi_{i},Y}Z,V)-\eta^{i}(Y)\,g(R_{\,\xi_{i},X}Z,V)]$
$+g(R_{{f}X,Y}{f}Z,V)-g(R_{{f}Y,X}{f}Z,V)+g(R_{{f}X,Y}Z,{f}V)-g(R_{{f}Y,X}Z,{f}V)$
$-g(R_{\widetilde{Q}X,Y}Z,V)-g(R_{X,\widetilde{Q}Y}Z,V)=0.$ (43)
Then, inserting $fZ$ in place of $Z$ and also $V$ by fV{f}V in (30) and using (2), we get two equations
$g(R_{X,Y}QZ,V)$ $=\sum\nolimits_{i}\eta^{i}(Z)\,g(R_{X,Y}\,\xi_{i},V)+g(R_{X,Y}{f}Z,{f}V)$
$\quad+g(R_{X,{f}Y}{f}Z,V)+g(R_{{f}X,Y}{f}Z,V),$ (44)
$g(R_{X,Y}Z,QV)$ $=\sum\nolimits_{i}\eta^{i}(V)\,g(R_{X,Y}Z,\,\xi_{i})+g(R_{X,Y}{f}Z,{f}V)$
$\quad+g(R_{{f}X,Y}Z,{f}V)+g(R_{X,{f}Y}Z,{f}V).$ (45)
Adding (4) to (4), and substituting the gotten equation into (4), we have
$2\,g(R_{{f}X,{f}Y}Z,V)-2\,g(R_{X,Y}{f}Z,{f}V)+\sum\nolimits_{i}\big[\eta^{i}(X)\,g(R_{\,\xi_{i},Y}Z,V)-\eta^{i}(Z)\,g(R_{X,Y}\,\xi_{i},V)$
$\quad-\eta^{i}(V)\,g(R_{X,Y}Z,\,\xi_{i})\,-\eta^{i}(Y)\,g(R_{\,\xi_{i},X}Z,V)\big]+\delta(X,Y,Z,V)=0,$ (46)
where
$\delta(X,Y,Z,V)=g(R_{X,Y}\widetilde{Q}Z,V)+g(R_{X,Y}Z,\widetilde{Q}V)-g(R_{\widetilde{Q}X,Y}Z,V)-g(R_{X,\widetilde{Q}Y}Z,V).$
Substituting $fX$ in place of $X$ and also $fY$ in place of $Y$ in (4) and using (2), we obtain
$2\,g(R_{QX,QY}Z,V)+2\,\sum\nolimits_{i}\big[\eta^{i}(Y)\,g(R_{\,\xi_{i},QX}Z,V)-\eta^{i}(X)\,g(R_{\,\xi_{i},QY}Z,V)\big]-2\,g(R_{{f}X,{f}Y}\,{f}Z,{f}V)$
$+\delta({f}X,{f}Y,Z,V)=\sum\nolimits_{i}\big[\eta^{i}(Z)\,g(R_{\,\xi_{i},V}{f}X,{f}Y)-\eta^{i}(V)\,g(R_{\,\xi_{i},Z}\,{f}X,{f}Y)\big].$ (47)
Inserting $X=\xi_{j}$ and $Y=\xi_{k}$ in (4), we acquire $g(R_{\xi_{j},\xi_{k}}Z,V)=0$ for all $1\leq j,k\leq s$; hence (31). This also yields $g(R_{\xi_{j},\,\xi_{k}}fZ,fV)=0$. Taking $V=\xi_{j}$ in (4) and using the previous relation, we obtain
$2\,g(R_{\,QX,\,QY}Z,\xi_{j})+2\sum\nolimits_{i}\big[\eta^{i}(Y)\,g(R_{\,\xi_{i},QX}Z,\xi_{j})-\eta^{i}(X)\,g(R_{\,\xi_{i},QY}Z,\xi_{j})\big]$
$+\,g(R_{\,\xi_{j},Z}\,{f}X,{f}Y)+\delta({f}X,{f}Y,Z,\,\xi_{j})=0\quad(1\leq j\leq s).$ (48)
Replacing $X$ by fX{f}X and $Y$ by fY{f}Y in (4), we obtain
$4\,g(R_{\,QfX,\,QfY}Z,\xi_{j})-2\sum\nolimits_{i}\big[\eta^{i}(Y)\,g(R_{\,\xi_{i},QX}Z,\xi_{j})-\eta^{i}(X)\,g(R_{\,\xi_{i},QY}Z,\xi_{j})\big]$
$+2\,g(R_{\,\xi_{j},Z}\,{Q}X,{Q}Y)+2\,\delta({f}^{2}X,{f}^{2}Y,Z,\,\xi_{j})=0\quad(1\leq j\leq s).$
Adding this to (4) and using $Q={\rm Id}+\widetilde{Q}$, we get
$3\,g(R_{\,\xi_{j},Z}\,{f}X,{f}Y)=-4\,g(R_{\,\xi_{j},Z}\,\widetilde{Q}{f}X,\,\widetilde{Q}{f}Y)-4\,g(R_{\,\xi_{j},Z}\,{f}X,\,\widetilde{Q}{f}Y)$
$-4\,g(R_{\,\xi_{j},Z}\,\widetilde{Q}{f}X,\,{f}Y)+\delta({f}X,{f}Y,Z,\,\xi_{j})+2\,\delta({f}^{2}X,{f}^{2}Y,Z,\,\xi_{j})\quad(1\leq j\leq s).$ (49)
From (4), using (13), (31) and the equality $\delta({f}X,{f}Y,Z,\xi_{j})=0$, we get (32). Since ${f}|_{\mathcal{D}}$ is non-degenerate, the distribution $\mathcal{D}$ is curvature invariant, see (1). ∎
Lemma 5.
For a weak nearly $\cal C$}-manifold satisfying conditions (6), (7), (10) and (13), we get
$R_{\,\xi_{i},\,X}Y=-(\nabla_{X}h_{i})Y,$ (50)
$(\nabla_{X}h_{i})Y=\sum\nolimits_{j}[g(h_{i}h_{j}X,Y)\,\xi_{j}-\eta^{j}(Y)\,h_{i}h_{j}X],$ (51)
${\rm Ric}\,(\xi_{i},Z)=-\sum\nolimits_{j}\eta^{j}(Z)\,\operatorname{tr}(h_{i}h_{j}).$ (52)
In particular, $\nabla_{\xi_{j}}h_{i}=0$ and ${\rm tr}(h_{i}^{2})=const$ for all $1\leq i,j\leq s$. Moreover,
$g(R_{\,\xi_{i},X}Y,Z)$ $=-g((\nabla_{X}\,h_{i})Y,Z)=\sum\nolimits_{j}[\eta^{j}(Y)\,g(h_{i}h_{j}X,Z)-\eta^{j}(Z)\,g(h_{i}h_{j}X,Y)],$ (53)
$g(R_{{f}X,{f}Y}Z,V)$ $=g(R_{X,Y}{f}Z,{f}V)-\tfrac{1}{2}\,\delta(X,Y,Z,V),$ (54)
$g(R_{{f}X,{f}Y}{f}Z,{f}V)$ $=g(R_{QX,QY}Z,V)+\tfrac{1}{2}\,\delta({f}X,{f}Y,Z,V)$
$\quad+\sum\nolimits_{i}\big[\eta^{i}(Y)\,g(R_{\,\xi_{i},QX}Z,V)-\eta^{i}(X)\,g(R_{\,\xi_{i},QY}Z,V)\big].$ (55)
Proof.
By Lemma 1, $\xi_{i}$ is a Killing vector field, and applying (12) in conjunction with (14), we get (50). Replacing $Y$ by fY{f}Y and $Z$ by fZ{f}Z in $g(R_{\,\xi_{i},X}Y,Z)=-g((\nabla_{X}\,h_{i})Y,Z)$, see (50), and using (32), we get $g((\nabla_{X}h_{i})\,{f}Y,{f}Z)=0$, hence,
[…]
Then, using (56), we find the $\ker f$-component and $\cal D$}-component of $(\nabla_{X}h_{i})Y$:
$g((\nabla_{X}h_{i})Y,\,\xi_{j})=g(\nabla_{X}(h_{i}\,Y),\,\xi_{j})=-g(h_{i}\,Y,\,\nabla_{X}\,\xi_{j})=g(h_{i}h_{j}X,Y),$
$g((\nabla_{X}\,h_{i})Y,Z)=\sum\nolimits_{j}\eta^{j}(Y)\,g((\nabla_{X}h_{i})\,\xi_{j},\,Z)=-\sum\nolimits_{j}\eta^{j}(Y)\,g(h_{i}h_{j}X,Z)\quad(Z\in{\cal D}),$
from which (51) follows. From (51) with $X=\xi_{k}$ we find $\nabla_{\xi_{k}}h_{i}=0$. Note that for $Y=\xi_{j}$, (51) reduces to (23). By (51) and (50), we get (53). Let $\{e_{k}\}_{k=1,\ldots,2n+1}$ be a local orthonormal frame on $M$ with $e_{2n+j}=\xi_{j}\ (1\leq j\leq s)$. Putting $X=Y=e_{k}$ in (51) for $\{e_{k}\}$, then using (1) and summing over $k=1,\ldots,2n+s$, we get
$\sum\nolimits_{k=1}^{2n+s}(\nabla_{e_{k}}h_{i})\,e_{k}=\sum\nolimits_{j}\operatorname{tr}(h_{i}h_{j})\xi_{j}.$
Next, taking the inner product of (50) with $Z$, tracing over $X$ and $Y$, and then substituting the above relation, we obtain (52). Replacing $Y$ by $h_{i}Y$ in (51), putting $Y=e_{k}$ in the gotten equation and summing over $k=1,\ldots,2n+s$, we get
$\operatorname{tr}((\nabla_{X}h_{i})h_{i})=\sum\nolimits_{k=1}^{2n+s}g((\nabla_{X}h_{i})h_{i}e_{k},e_{k})=\sum\nolimits_{j}\sum\nolimits_{k=1}^{2n+s}\eta^{j}(e_{k})\,g(h_{i}h_{j}X,h_{i}e_{k})=0.$
This implies $X({\rm tr}(h_{i}^{2}))=0\ (X\in\mathfrak{X}_{M})$, i.e., ${\rm tr}(h_{i}^{2})=const$. Finally, using (4) and (53), we obtain
$2g(R_{{f}X,{f}Y}Z,V)-2g(R_{X,Y}{f}Z,{f}V)+\!\sum\nolimits_{i,j=1}^{s}\!\big[\eta^{i}(X)\eta^{j}(Z)g(h_{i}h_{j}Y,V)-\eta^{i}(Z)\eta^{j}(X)g(h_{i}h_{j}V,Y)$
$-\eta^{i}(X)\eta^{j}(V)g(h_{i}h_{j}Y,Z)+\eta^{i}(V)\eta^{j}(X)g(h_{i}h_{j}Z,Y)+\eta^{i}(Z)\eta^{j}(Y)g(h_{i}h_{j}V,X)$
$-\eta^{i}(Y)\eta^{j}(Z)g(h_{i}h_{j}X,V){+}\eta^{i}(Y)\eta^{j}(V)g(h_{i}h_{j}X,Z){-}\eta^{i}(V)\eta^{j}(Y)g(h_{i}h_{j}Z,X)\big]+\delta(X,Y,Z,V)=0.$
Simplification of this using the skew-symmetry of $h_{i}$ yields (54). Switching $X$ by fX{f}X and $Y$ by fY{f}Y in (54) and using (2), we get (5). ∎
Remark 6.
The function $\delta$ of a weak nearly $\cal C$}-manifold has the following symmetries:
$\delta(Y,X,Z,V)=\delta(X,Y,V,Z)=\delta(Z,V,X,Y)=-\delta(X,Y,Z,V).$
If (13) is true, then by (53), we acquire
$\delta(\xi_{i},Y,Z,V)\!=\delta(X,\xi_{i},Z,V)\!=\delta(X,Y,\xi_{i},V)\!=\delta(X,Y,Z,\xi_{i})=0\quad(1\leq i\leq s).$
Lemma 6.
For a weak nearly $\cal C$}-manifold satisfying (6), (7), (10) and (13), we have for $1\leq j\leq s$
$g((\nabla_{X}{f})Y,Q{f}h_{j}Z){=}\sum\nolimits_{i}\big[\eta^{i}(X)g(h_{j}Y,h_{i}Z){-}\eta^{i}(Y)g(h_{j}X,h_{i}Z){+}\eta^{i}(X)g(\widetilde{Q}(Q{+}{\rm Id})Y,h_{i}h_{j}Z)\big].$ (57)
Proof.
It consists of two parts. In Part 1, our formulas depend on four vectors from $TM$ and contain many additional $\widetilde{Q}-$dependent terms. In Part 2, the formulas depend on three vectors from $TM$ and contain a few $\widetilde{Q}-$dependent terms.
Part 1. Differentiating (27) and using $g((\nabla_{X}{f})(\nabla_{V}{f})Y,Z)$ $g((\nabla_{V}{f})Y,(\nabla_{X}{f})Z)$ gives
$g((\nabla_{V}\,{f})Y,(\nabla_{X}{f})Z)+g((\nabla_{X}{f})Y,(\nabla_{V}\,{f})Z)=g((\nabla^{2}_{V,X}\,{f}){f}Y,Z)+g((\nabla^{2}_{V,X}\,{f}){f}Z,Y)$
$-\sum\nolimits_{i}\big[g(h_{i}V,Z)\,g(h_{i}X,Y)+\eta^{i}(Z)\,g((\nabla_{V}h_{i})X,Y)+g(h_{i}V,Y)\,g(h_{i}X,Z)$
$+\eta^{i}(Y)\,g((\nabla_{V}h_{i})X,Z)+\eta^{i}(Y)g((\nabla_{V}h_{i})X,\widetilde{Q}Z)+\eta^{i}(Z)g((\nabla_{V}h_{i})X,\widetilde{Q}Y)$
$+g(h_{i}V,Y)g(h_{i}X,\widetilde{Q}Z)+g(h_{i}V,Z)g(h_{i}X,\widetilde{Q}Y)\big]+2\sum\nolimits_{i,j}\eta^{i}(Y)\eta^{j}(Z)g(h_{i}X,\widetilde{Q}h_{j}V).$ (58)
Using (2), (33) and (39), we find $\nabla^{2}$-terms in (4):
$g((\nabla^{2}_{V,X}\,{f}){f}Z,Y)=g((\nabla^{2}_{V,{f}Z}\,{f})Y,X)=-g(R_{X,Y}V,QZ)+\sum\nolimits_{i}\eta^{i}(Z)\,g(R_{X,Y}V,\xi_{i})$
$\qquad-\tfrac{1}{2}\,g(R_{X,V}{f}Z,{f}Y)-\tfrac{1}{2}\,g(R_{V,Y}{f}Z,{f}X)-\tfrac{1}{2}\,g(R_{X,Y}{f}Z,{f}V),$ (59)
$g((\nabla^{2}_{V,X}\,{f}){f}Y,Z)=g((\nabla^{2}_{V,{f}Y}\,{f})Z,X)=-g(R_{X,Z}V,QY)+\sum\nolimits_{i}\eta^{i}(Y)\,g(R_{X,Z}V,\xi_{i})$
$\qquad-\tfrac{1}{2}\,g(R_{X,V}{f}Y,{f}Z)-\tfrac{1}{2}\,g(R_{V,Z}{f}Y,{f}X)-\tfrac{1}{2}\,g(R_{X,Z}{f}Y,{f}V).$ (60)
Substituting (4)–(4) in (4), we get
$g((\nabla_{V}\,{f})Y,(\nabla_{X}{f})Z)+g((\nabla_{X}{f})Y,(\nabla_{V}\,{f})Z)=-g(R_{X,Z}V,QY)-g(R_{X,Y}V,QZ)$
$-\tfrac{1}{2}\,g(R_{V,Z}{f}Y,{f}X)-\tfrac{1}{2}\,g(R_{X,Z}{f}Y,{f}V)-\tfrac{1}{2}\,g(R_{V,Y}{f}Z,{f}X)-\tfrac{1}{2}\,g(R_{X,Y}{f}Z,{f}V)$
$+\sum\nolimits_{i}\big[\eta^{i}(Y)\,g(R_{X,Z}V,\xi_{i})+\eta^{i}(Z)\,g(R_{X,Y}V,\xi_{i})-g(h_{i}V,Z)\,g(h_{i}X,Y)-\eta^{i}(Z)\,g((\nabla_{V}h_{i})X,Y)$
$-g(h_{i}V,Y)\,g(h_{i}X,Z)-\eta^{i}(Y)\,g((\nabla_{V}h_{i})X,Z)-\eta^{i}(Y)g((\nabla_{V}h_{i})X,\widetilde{Q}Z)-\eta^{i}(Z)g((\nabla_{V}h_{i})X,\widetilde{Q}Y)$
$-g(h_{i}V,Y)g(h_{i}X,\widetilde{Q}Z)-g(h_{i}V,Z)g(h_{i}X,\widetilde{Q}Y)\big]+2\sum\nolimits_{i,j}\eta^{i}(Y)\eta^{j}(Z)g(h_{i}X,\widetilde{Q}h_{j}V).$ (61)
Applying (54) twice followed by (51) in (4), the curvature terms involving $\xi_{i}$ vanish, and we get
$g((\nabla_{X}{f})Z,(\nabla_{V}{f})Y)+g((\nabla_{X}{f})Y,(\nabla_{V}{f})Z)+g(R_{X,Z}V,QY)+g(R_{X,Y}V,QZ)$
$-g(R_{V,Z}{f}X,{f}Y)-g(R_{X,Z}{f}V,{f}Y)+\sum\nolimits_{i}\big[g(h_{i}V,Y)\,g(h_{i}X,Z)+g(h_{i}V,Z)\,g(h_{i}X,Y)\big]$
$=$ $-\tfrac{1}{4}\delta(X,Z,V,Y){-}\tfrac{1}{4}\delta(V,Z,X,Y){-}\sum\nolimits_{i}[g(h_{i}V,Y)g(h_{i}X,\widetilde{Q}Z)+g(h_{i}V,Z)g(h_{i}X,\widetilde{Q}Y)\big]$
$+\sum\nolimits_{i,j}\big[2\eta^{i}(Y)\eta^{j}(Z)g(h_{i}X,\widetilde{Q}h_{j}V){-}\eta^{i}(Y)\eta^{j}(X)g(h_{i}h_{j}V,\widetilde{Q}Z){-}\eta^{i}(Z)\eta^{j}(X)g(h_{i}h_{j}V,\widetilde{Q}Y)\big].$ (62)
Replacing $Z$ by fZ{f}Z and $V$ by fV{f}V in (4), we find
$g((\nabla_{X}{f}){f}Z,(\nabla_{{f}V}\,{f})Y)+g((\nabla_{X}{f})Y,(\nabla_{{f}V}\,{f}){f}Z)+g(R_{X,{f}Z}{f}V,QY){-}g(R_{X,{f}Z}{f}^{2}V,{f}Y)$
$+g(R_{X,Y}{f}V,{f}QZ)-g(R_{{f}V,{f}Z}{f}X,{f}Y)+\sum\nolimits_{i}\big[g(h_{i}X,Y)\,g(h_{i}{f}V,{f}Z)+g(h_{i}X,{f}Z)\,g(h_{i}{f}V,Y)\big]$
$=-\tfrac{1}{4}\,\delta(X,{f}Z,{f}V,Y)-\tfrac{1}{4}\,\delta({f}V,{f}Z,X,Y)-\sum\nolimits_{i,j}\eta^{i}(Y)\eta^{j}(X)g(h_{i}h_{j}fV,\widetilde{Q}fZ)$
$-\sum\nolimits_{i}[g(h_{i}fV,Y)g(h_{i}X,\widetilde{Q}fZ)+g(h_{i}fV,fZ)g(h_{i}X,\widetilde{Q}Y)\big].$ (63)
Using (2), (27), (28), (3) and Lemma 2, we find
$g((\nabla_{X}{f}){f}Z,(\nabla_{{f}V}\,{f})Y)=g(Q(\nabla_{X}{f})Z,(\nabla_{V}\,{f})Y)-\sum\nolimits_{i,j}\eta^{i}(Z)\,\eta^{j}(V)g(Qh_{i}X,Qh_{j}Y)$
$\quad+\sum\nolimits_{i}\big[\eta^{i}(Z)g(Q{f}h_{i}X,(\nabla_{V}\,{f})Y)-g(X,{f}h_{i}Z)\,g(V,{f}h_{i}Y)-\eta^{i}(V)\,g((\nabla_{X}{f})Z,Q{f}h_{i}Y)$
$\quad+g(h_{i}X,QZ)g(h_{i}V,QY)\big],$ (64)
$g((\nabla_{X}{f})Y,(\nabla_{{f}V}{f}){f}Z)=-g(Q(\nabla_{X}{f})Y,(\nabla_{V}\,{f})Z)+\sum\nolimits_{i}\big[\eta^{i}(V)g(Q{f}h_{i}Z,(\nabla_{X}{f})Y)$
$\quad-\eta^{i}(Z)g(Q{f}h_{i}V,(\nabla_{X}{f})Y)+g({f}h_{i}X,Y)g({f}h_{i}V,\widetilde{Q}Z)\big].$ (65)
From (53) and Lemma 2, we have
$g(R_{X,{f}Z}{f}V,Y)-g(R_{X,{f}Z}{f}^{2}V,{f}Y)$ $=g(R_{X,{f}Z}{f}V,Y)+g(R_{X,{f}Z}QV,{f}Y)$
$\quad-\sum\nolimits_{i,j}\eta^{i}(X)\,\eta^{j}(V)\,g(h_{j}h_{i}Z,QY).$ (66)
On the other hand, from (30), (54) and (2) it follows that
$g(R_{X,Z}{f}V,{f}Y)+g(R_{X,Z}{f}^{2}V,Y)+g(R_{X,{f}Z}{f}V,Y)+g(R_{{f}X,Z}{f}V,Y)=0,$ (67)
- $g(R_{{f}X,Z}{f}V,QY)+\sum\nolimits_{i}\eta^{i}(Y)\,g(R_{{f}X,Z}{f}V,\xi_{i})=-g(R_{QX,{f}Z}V,{f}Y)$
$\quad+\sum\nolimits_{i}\eta^{i}(X)\,g(R_{\xi_{i}\,,{f}Z}V,{f}Y)+\tfrac{1}{2}\,\delta({f}X,Z,V,{f}Y).$ (68)
Summing up the formulas (67) and (4) (and using (2), (53) and Lemma 2), we obtain
$g(R_{X,{f}Z}{f}V,Y)$ $+g(R_{QX,{f}Z}V,{f}Y)=g(R_{X,Z}QV,Y)-g(R_{X,Z}{f}V,{f}Y)+g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)$
$+\sum\nolimits_{i,j}\big[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)-\eta^{i}(Y)\,\eta^{j}(Z)\,g(h_{i}h_{j}V,QX)$
$+\eta^{i}(Z)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,X)\big]+\tfrac{1}{2}\,\delta({f}X,Z,V,{f}Y).$ (69)
Substituting (4) into (4), we get
$g(R_{X,{f}Z}\,{f}V,Y)-g(R_{X,{f}Z}\,{f}^{2}V,{f}Y)=g(R_{X,Z}\,QV,Y)-g(R_{X,Z}{f}V,{f}Y)$
$+\sum\nolimits_{i,j=1}^{s}\big[\eta^{i}(Z)\eta^{j}(V)\,g(h_{i}h_{j}Y,X)-\eta^{i}(Y)\eta^{j}(Z)\,g(h_{j}h_{i}X,QV)-\eta^{i}(X)\,\eta^{j}(V)\,g(h_{j}h_{i}Z,Y)\big]$
$-g(R_{\widetilde{Q}X,{f}Z}V,{f}Y)+g(R_{X,{f}Z}\widetilde{Q}V,{f}Y)+g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)+\tfrac{1}{2}\,\delta({f}X,Z,V,{f}Y).$ (70)
Using (53) and (54)–(5), we have
$g(R_{X,Y}{f}V,{f}Z)-g(R_{{f}V,{f}Z}{f}X,{f}Y)=g(R_{V,Z}{f}X,{f}Y)-g(R_{V,Z}QX,QY)$
$+\sum\nolimits_{i,j}\big[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}QY,Z)-\eta^{i}(X)\eta^{j}(Z)\,g(h_{i}h_{j}QY,V)+\,\eta^{i}(Y)\,\eta^{j}(Z)\,g(h_{i}h_{j}QX,V)$
$-\eta^{i}(Y)\,\eta^{j}(V)\,g(h_{i}h_{j}QX,Z)\big]-\tfrac{1}{2}\,\delta({f}X,{f}Y,Z,V)-\tfrac{1}{2}\,\delta(X,Y,Z,V).$ (71)
Applying (4), (4), (4) and (4) in (4), and using Lemma 2, we obtain
$g(Q(\nabla_{X}{f})Z,(\nabla_{V}{f})Y)-g(Q(\nabla_{X}{f})Y,(\nabla_{V}{f})Z)+\sum\nolimits_{i}\big[\eta^{i}(V)g((\nabla_{X}{f})Y,Q{f}h_{i}Z)$
$-\eta^{i}(V)\,g((\nabla_{X}{f})Z,Q{f}h_{i}Y)+\eta^{i}(Z)g((\nabla_{V}{f})Y,Q{f}h_{i}X)-\eta^{i}(Z)g((\nabla_{X}{f})Y,Q{f}h_{i}V)$
$+g(h_{i}X,QZ)g(h_{i}V,QY)-g(h_{i}X,Y)\,g(h_{i}V,QZ)\big]+\sum\nolimits_{i,j}\big[2\,\eta^{i}(Z)\eta^{j}(V)g(h_{i}h_{j}Y,X)$
$-\eta^{i}(X)\,\eta^{j}(Z)\,g(h_{i}h_{j}Y,QV)-\eta^{i}(Y)\,\eta^{j}(V)\,g(h_{i}h_{j}X,QZ)\big]+g(R_{X,Z}\,QV,Y)-g(R_{X,Z}{f}V,{f}Y)$
$+g(R_{V,Z}{f}X,{f}Y)-g(R_{V,Z}\,QX,QY)=g(R_{\widetilde{Q}X,{f}Z}V,{f}Y)-g(R_{X,{f}Z}\widetilde{Q}V,{f}Y)$
$-g(R_{X,{f}Z}{f}V,\widetilde{Q}Y)-g(R_{X,Y}{f}V,{f}\widetilde{Q}Z)-g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)-\sum\nolimits_{i}[g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)$
$+g(h_{i}fV,Y)g(h_{i}X,\widetilde{Q}fZ)+g(h_{i}fV,fZ)g(h_{i}X,\widetilde{Q}Y)\big]+\sum\nolimits_{i,j}^{s}[\eta^{i}(Z)\eta^{j}(V)g(\widetilde{Q}(Q{+}{\rm Id})h_{i}X,h_{j}Y)$
$-\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)-\eta^{i}(Y)\eta^{j}(X)g(h_{i}h_{j}fV,\widetilde{Q}fZ)]+\tfrac{1}{2}\,\delta({f}X,{f}Y,Z,V)$
$-\tfrac{1}{2}\,\delta({f}X,Z,V,{f}Y)+\tfrac{1}{2}\,\delta(X,Y,Z,V)-\tfrac{1}{4}\,\delta(X,{f}Z,{f}V,Y)-\tfrac{1}{4}\,\delta({f}V,{f}Z,X,Y).$ (72)
Adding (4) to (4), we obtain
$2\,g((\nabla_{X}{f})Z,(\nabla_{V}{f})Y)+g(\widetilde{Q}(\nabla_{X}{f})Z,(\nabla_{V}{f})Y)-g(\widetilde{Q}(\nabla_{X}{f})Y,(\nabla_{V}{f})Z)$
$+\sum\nolimits_{i}\big[\eta^{i}(V)g((\nabla_{X}{f})Y,Q{f}h_{i}Z)-\eta^{i}(V)\,g((\nabla_{X}{f})Z,Q{f}h_{i}Y)+\eta^{i}(Z)g((\nabla_{V}{f})Y,Q{f}h_{i}X)$
$-\eta^{i}(Z)g((\nabla_{X}{f})Y,Q{f}h_{i}V)+g(h_{i}X,QZ)g(h_{i}V,QY)+g(h_{i}V,Y)\,g(h_{i}X,Z)\big]$
$+\sum\nolimits_{i,j}^{s}\big[2\,\eta^{i}(Z)\eta^{j}(V)g(h_{i}h_{j}Y,X)-\eta^{i}(X)\eta^{j}(Z)g(h_{i}h_{j}Y,QV)-\eta^{i}(Y)\eta^{j}(V)g(h_{i}h_{j}X,QZ)\big]$
$+2\,g(R_{X,Z}\,V,Y)-2\,g(R_{X,Z}{f}V,{f}Y)-g(R_{V,Z}\,QX,QY)+g(R_{X,Y}V,QZ)$
$=g(R_{\widetilde{Q}X,{f}Z}V,{f}Y){-}g(R_{X,{f}Z}\widetilde{Q}V,{f}Y){-}g(R_{X,{f}Z}{f}V,\widetilde{Q}Y)-g(R_{X,Z}\,\widetilde{Q}V,Y)-g(R_{X,Z}V,\widetilde{Q}Y)$
$-g(R_{X,Y}{f}V,{f}\widetilde{Q}Z)-g(R_{{f}X,Z}{f}V,\widetilde{Q}Y)-\sum\nolimits_{i,j}[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Y,\widetilde{Q}Z)$
$-\eta^{i}(Z)\eta^{j}(V)g(\widetilde{Q}(Q{+}{\rm Id})h_{i}X,h_{j}Y)-2\eta^{i}(Y)\eta^{j}(Z)g(h_{i}X,\widetilde{Q}h_{j}V)+\eta^{i}(Y)\eta^{j}(X)g(h_{i}h_{j}V,\widetilde{Q}Z)$
$+\eta^{i}(Z)\eta^{j}(X)g(h_{i}h_{j}V,\widetilde{Q}Y)+\eta^{i}(Y)\eta^{j}(X)g(h_{i}h_{j}fV,\widetilde{Q}fZ)]+\sum\nolimits_{i}\big[g(h_{i}X,Y)\,g(h_{i}V,\widetilde{Q}Z)$
$-g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V){-}g(h_{i}fV,Y)g(h_{i}X,\widetilde{Q}fZ){-}g(h_{i}fV,fZ)g(h_{i}X,\widetilde{Q}Y){-}g(h_{i}V,Y)g(h_{i}X,\widetilde{Q}Z)$
$-g(h_{i}V,Z)g(h_{i}X,\widetilde{Q}Y)]-\tfrac{1}{2}\,\delta({f}X,Z,V,{f}Y)+\tfrac{1}{2}\,\delta({f}X,{f}Y,Z,V)+\tfrac{1}{2}\,\delta(X,Y,Z,V)$
$-\tfrac{1}{4}\delta(X,Z,V,Y)-\tfrac{1}{4}\delta(V,Z,X,Y)-\tfrac{1}{4}\,\delta(X,{f}Z,{f}V,Y)-\tfrac{1}{4}\,\delta({f}V,{f}Z,X,Y).$ (73)
Swapping $X\leftrightarrow Z$ and $V\leftrightarrow Y$ in (4), then subtracting the gotten equation from (4) and using (9), we get
$\sum\nolimits_{i}\big[\eta^{i}(V)g((\nabla_{X}{f})Y,Q{f}h_{i}Z)-\eta^{i}(V)\,g((\nabla_{X}{f})Z,Q{f}h_{i}Y)+\eta^{i}(Z)g((\nabla_{V}\,{f})Y,Q{f}h_{i}X)$
$-\eta^{i}(Y)\,g((\nabla_{Z}\,{f})V,Q{f}h_{i}X)-\eta^{i}(Z)g((\nabla_{X}{f})Y,Q{f}h_{i}V)+\eta^{i}(Y)g((\nabla_{Z}{f})X,Q{f}h_{i}V)$
$-\eta^{i}(X)g((\nabla_{Y}\,{f})V,Q{f}h_{i}Z)+\eta^{i}(X)g((\nabla_{Z}{f})V,Q{f}h_{i}Y)\big]+g(R_{Z,V}\,QX,QY)+g(R_{Z,V}\,QX,Y)$
$-g(R_{X,Y}\,QZ,V){-}g(R_{X,Y}\,QZ,QV){+}2\sum\nolimits_{i,j}\big[\eta^{i}(Z)\eta^{j}(V)g(h_{i}h_{j}Y,X){-}\eta^{i}(X)\eta^{j}(Y)g(h_{i}h_{j}V,Z)]$
$=g(R_{\widetilde{Q}X,{f}Z}V,{f}Y){-}g(R_{X,{f}Z}{f}V,\widetilde{Q}Y)+g(R_{Z,V}{f}Y,{f}\widetilde{Q}X){-}g(R_{X,Y}{f}V,{f}\widetilde{Q}Z){-}g(R_{\widetilde{Q}Z,{f}X}Y,{f}V)$
$+g(R_{Z,{f}X}{f}Y,\widetilde{Q}V)+\sum\nolimits_{i,j}\big[\eta^{i}(X)\,\eta^{j}(V)\,g(h_{i}h_{j}Z,\widetilde{Q}Y)-\eta^{i}(Z)\,\eta^{j}(Y)\,g(h_{i}h_{j}V,\widetilde{Q}X)\big]$
$+\sum\nolimits_{i}\big[g(h_{i}X,Y)\,g(h_{i}V,\widetilde{Q}Z)-g({f}h_{i}X,Y)g({f}h_{i}Z,\widetilde{Q}V)-g(h_{i}V,QZ)g(h_{i}X,\widetilde{Q}Y)$
$-g(h_{i}V,Z)g(h_{i}X,\widetilde{Q}Y){-}g(h_{i}Z,V)\,g(h_{i}Y,\widetilde{Q}X){+}g({f}h_{i}Z,V)g({f}h_{i}X,\widetilde{Q}Y){+}g(h_{i}Y,QX)g(h_{i}Z,\widetilde{Q}V)$
$-g(h_{i}Y,X)g(h_{i}Z,\widetilde{Q}V)\big]+\tfrac{1}{4}\delta({f}X,Z,{f}Y,V)+\tfrac{1}{4}\,\delta({f}X,{f}Y,Z,V)+\tfrac{1}{2}\,\delta(X,Y,Z,V)$
$+\tfrac{1}{4}\,\delta(X,{f}Z,{f}V,Y)-\tfrac{1}{4}\,\delta({f}Z,{f}V,X,Y).$ (74)
Part 2. Inserting $\xi_{k}$ in place of $V$ in (4), then using Lemma 2, (1), Remark 6 and (9), we get
$g((\nabla_{X}{f})Y,Q{f}h_{k}Z)=g((\nabla_{X}{f})Z,Q{f}h_{k}Y){-}\sum\nolimits_{i}\big[\eta^{i}(Z)g(Qh_{k}Y,Qh_{i}X){+}\eta^{i}(Y)g(Qh_{k}Z,Qh_{i}X)$
$+2\eta^{i}(Z)g(h_{k}Y,h_{i}X){-}\eta^{i}(X)g(h_{i}h_{k}Y,\widetilde{Q}Z)\big]{+}g(R_{\xi_{k},Z}QX,QY){+}g(R_{\xi_{k},Z}QX,Y){-}2g(R_{\xi_{k},QZ}X,Y).$
Simplifying the above equation using (2) and (53), we have
$g((\nabla_{X}{f})Y,Q{f}h_{k}Z)$ $=g((\nabla_{X}{f})Z,Q{f}h_{k}Y){+}\sum\nolimits_{i}\big[\eta^{i}(Z)g(Qh_{k}Y,Qh_{i}X){-}\eta^{i}(Y)g(Qh_{k}Z,Qh_{i}Y)$
$\quad+2\,\eta^{i}(Z)g(\widetilde{Q}(Q+{\rm Id})X,h_{i}h_{k}Y)\big].$ (75)
By swapping $X$ and $Y$ in (4) and adding this to (4), we get
$g((\nabla_{Y}f)Z,Qfh_{k}X)+g((\nabla_{X}f)Z,Qfh_{k}Y)=\sum\nolimits_{i}\big[\eta^{i}(X)g(Qh_{i}Y,Qh_{k}Z)+\eta^{i}(Y)g(Qh_{i}X,Qh_{k}Z)$
$-2\eta^{i}(Z)g(Qh_{i}X,Qh_{k}Y)+4\eta^{i}(X)g(\widetilde{Q}(Q+{\rm Id})Y,h_{i}h_{k}Z)\big].$ (76)
By swapping $Z$ and $X$ in (4), and adding this to (4), we get
$g((\nabla_{X}f)Y,Qfh_{k}Z)=\sum\nolimits_{i}\big[\eta^{i}(X)g(Qh_{i}Z,Qh_{k}Y)-\eta^{i}(Z)g(Qh_{k}Y,Qh_{i}X)$
$\quad+2\eta^{i}(X)g(\widetilde{Q}(Q+{\rm Id})X,h_{i}h_{k}Z)-\eta^{i}(Z)g(\widetilde{Q}(Q+{\rm Id})X,h_{i}h_{k}Y)\big].$
Finally, simplifying the above equation using the definition of $\widetilde{Q}$, we get the desired (57). ∎
Next, we generalize Proposition 3.2 in [16] to the case $s>1$.
Lemma 7.
Let $\Omega^{k}(M)$ be the space of $k$-forms on an $s$-contact manifold $(M^{2n+s},\vec{\eta},{\mathcal{R}}\oplus{\cal D})$. If $n\geq 3$, then the map $\mathcal{E}_{d\eta^{i}}:\ \Omega^{2}(M)\longrightarrow\Omega^{4}(M),\,\beta\mapsto d\eta^{i}\wedge\beta$ is injective for all $i=1,\ldots,s$.
Proof.
We denote by $\Omega^{k}(\mathcal{D})$ the space of $k$-forms along $\mathcal{D}$, i.e., $k$-forms vanishing whenever at least one argument lies in $\mathcal{R}$}. The proof of the proposition consists of four steps.
Step 1 (Symplecticity on $\mathcal{D}$ and Lefschetz injectivity): By the definition of an $s$-contact manifold, $d\eta^{i}|_{\mathcal{D}}$ is nondegenerate at every point $p\in M$. Each fibre $(\mathcal{D}_{p},d\eta^{i}|_{{\mathcal{D}}_{p}})$ is a $2n$-dimensional symplectic vector space, hence $(d\eta^{i}|_{{\mathcal{D}}})^{n}$ is nowhere vanishing. In particular, by [5, Proposition 1.1] (the algebraic Lefschetz property), the linear maps $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\,\Omega^{2}(\mathcal{D})\longrightarrow\Omega^{4}(\mathcal{D})$, $\mathcal{E}_{d\eta^{i}|_{{\mathcal{D}}}}:\,\Omega^{1}(\mathcal{D})\longrightarrow\Omega^{3}(\mathcal{D})$ are injective for $n\geq 3$ and $n\geq 2$, respectively (the latter injectivity, in particular, holds when $n\geq 3$).
Step 2 (Bundle-level decomposition by $\eta$-degree): At each $p\in M$ we have a direct-sum decomposition $T_{p}^{*}M={\mathrm{span}\{\eta^{1}_{p},\dots,\eta^{s}_{p}\}}\ \oplus\ \mathcal{D}_{p}^{*},$ which induces a decomposition of exterior powers: $\Lambda^{k}(T_{p}^{*}M)\;=\;\bigoplus\nolimits_{j=0}^{\min(k,s)}\Lambda^{j}(\mathrm{span}\{\eta^{1}_{p},\dots,\eta^{s}_{p}\})\,\wedge\,\Lambda^{k-j}\mathcal{D}_{p}^{*}$. Since the ranks are constant and the summands depend smoothly on $p$, this yields a decomposition of bundles and spaces of sections:
$\Omega^{k}(M)\;=\;\bigoplus\nolimits_{j=0}^{\min(k,s)}{\Lambda^{j}\!\big(\mathrm{span}\{\eta^{1},\dots,\eta^{s}\}\big)}\wedge\Omega^{k-j}(\mathcal{D}).$
In particular, every $\beta\in\Omega^{2}(M)$ admits a unique splitting
$\beta\;=\;\beta_{0}\;+\;\sum\nolimits_{a=1}^{s}\eta^{a}\wedge\mu_{a}\;+\;\sum\nolimits_{1\leq a<b\leq s}f_{ab}\,\eta^{a}\wedge\eta^{b},$ (77)
with $\beta_{0}\in\Omega^{2}(\mathcal{D})$, $\mu_{a}\in\Omega^{1}(\mathcal{D})$ and $f_{ab}\in C^{\infty}(M)$.
Step 3 ($d\eta^{i}$ has no $\eta$-components): Expand $d\eta^{i}$ according to the same decomposition:
$d\eta^{i}\;=\;d\eta^{i}|_{\mathcal{D}}+\sum\nolimits_{a=1}^{s}\eta^{a}\wedge\gamma_{a}\;+\;\sum\nolimits_{1\leq a<b\leq s}h_{ab}\,\eta^{a}\wedge\eta^{b},$
with $\gamma_{a}\in\Omega^{1}(\mathcal{D})$ and $h_{ab}\in C^{\infty}(M)$. Let $\{\xi_{1},\dots,\xi_{s}\}$ be a local frame of $\mathcal{R}$} dual to $\{\eta^{1},\dots,\eta^{s}\}$. Using $\ker\,d\eta^{i}={\mathcal{R}}$, we have $\iota_{\xi_{a}}d\eta^{i}=0$ for each $i$ and $a$. Contracting the above expansion yields
$0=\iota_{\xi_{a}}d\eta^{i}=\ \gamma_{a}\;+\;\sum\nolimits_{b}h_{ab}\,\eta^{b}\;-\;\sum\nolimits_{b}h_{ba}\,\eta^{b}=\ \gamma_{a}\;+\;2\sum\nolimits_{b}h_{ab}\,\eta^{b}.$
Since $\{\eta^{b}\}$ are pointwise independent, it follows that $\gamma_{a}=0$ and $h_{ab}=0$ for all $a,b$. Hence $d\eta^{i}=d\eta^{i}|_{\mathcal{D}}$, that means that $d\eta^{i}$ has no terms involving $\eta^{i}$.
Step 4 (Wedge and compare $\eta$-degrees): Assume $d\eta^{i}\wedge\beta=0$. Using (77),
$0=d\eta^{i}|_{\mathcal{D}}\wedge\beta$ $={d\eta^{i}|_{\mathcal{D}}\wedge\beta_{0}}+\sum\nolimits_{a=1}^{s}{\eta^{a}\wedge\bigl(d\eta^{i}|_{\mathcal{D}}\wedge\mu_{a}\bigr)}+\big(\sum\nolimits_{1\leq a<b\leq s}{f_{ab}\,\eta^{a}\wedge\eta^{b}\big)\wedge d\eta^{i}|_{\mathcal{D}}}.$ (78)
Since different $\eta$-degrees lie in a direct sum, each component vanishes separately:
$\bullet$ Zero $\eta$-part (the first term of (78)): $d\eta^{i}|_{\mathcal{D}}\wedge\beta_{0}=0$ in $\Omega^{4}({\mathcal{D}})$. Now, as $d\eta^{i}|_{\mathcal{D}}\in\Omega^{2}({\mathcal{D}})$ and by the decomposition (77), $\beta_{0}\in\Omega^{2}({\mathcal{D}})$ and since $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\Omega^{2}({\mathcal{D}})\to\Omega^{4}({\mathcal{D}})$ is injective for $n\geq 3$ (by [5, Proposition 1.1]), we get $\beta_{0}=0$.
$\bullet$ One $\eta$-part: For each $a$, $d\eta^{i}|_{\mathcal{D}}\wedge\mu_{a}=0$ in $\Omega^{3}({\mathcal{D}})$. Again, since $\mathcal{E}_{d\eta^{i}|_{\mathcal{D}}}:\Omega^{1}({\mathcal{D}})\to\Omega^{3}({\mathcal{D}})$ is injective (in particular, for $n\geq 3$), $\mu_{a}=0$ for all $a$.
$\bullet$ Two $\eta$-parts: From Step 4 we are left with $\big(\sum\nolimits_{1\leq a<b\leq s}f_{ab}\,\eta^{a}\wedge\eta^{b}\big)\wedge d\eta^{i}|_{\mathcal{D}}=0$. Wedge this identity with $(d\eta^{i}|_{\mathcal{D}})^{\,n-1}$ to obtain $\big(\sum\nolimits_{a<b}f_{ab}\,\eta^{a}\wedge\eta^{b}\big)\wedge(d\eta^{i}|_{\mathcal{D}})^{\,n}=0$. By Step 1, $(d\eta^{i}|_{\mathcal{D}})^{\,n}$ is a nowhere-vanishing horizontal top form on $\mathcal{D}$}. Fix any point $p\in M$. Then $\{\,\eta^{a}_{p}\wedge\eta^{b}_{p}\,\}_{a<b}$ is a basis of $\Lambda^{2}(\mathrm{span}\{\eta^{1}_{p},\dots,\eta^{s}_{p}\})$, and wedging with $(d\eta^{i}|_{\mathcal{D}})^{\,n}\neq 0$ multiplies these basis elements by the same nonzero horizontal volume form. Hence, $\{\,\eta^{a}_{p}\wedge\eta^{b}_{p}\,\wedge(d\eta^{i}|_{\mathcal{D}})^{n}\}_{a<b}$ are linearly independent $(2n+2)$-forms on $M^{2n+s}$ where $s>1$. Applying this and the equality above forces $f_{ab}(p)=0$ for $a<b$. Since $p\in M$ is arbitrary, $f_{ab}\equiv 0$ on $M$. Thus, $\beta=0$, and $\mathcal{E}_{d\eta^{i}}$ is injective on $\Omega^{2}(M)$ for $n\geq 3$. ∎
Conclusion
We have shown that the weak nearly $\cal C$}-structure is useful for studying metric $f$-structures, almost contact metric structures, and Killing vector fields. Some results on nearly cosymplectic manifolds and weak nearly cosymplectic manifolds (see [10, 16, 20]) were generalized to weak nearly $\cal C$}-manifolds satisfying (10) and (13). Namely, the splitting Theorem 2 was proven and the $(4+s)$-dimensional case was characterized in Theorem 1. Their consequences (without the requirement (10) and (13)) present new results for nearly $\cal C$}-manifolds. Based on the numerous applications of the metric $f$-structure, we expect that the weak nearly $\cal C$}-structure will also be useful for differential geometry (e.g., the theory of $\mathfrak{g}$-foliations, $s$-cosymplectic structures and $s$-contact structures) and theoretical physics.
Statements and Declarations
  • Funding: Sourav Nayak is financially supported by a UGC research fellowship (Grant No. 211610029330). Dhriti Sundar Patra would like to thank the Science and Engineering Research Board (SERB), India, for financial support through the Start-up Research Grant (SRG) (Grant No. SRG/2023/002264).
  • Conflict of interest/Competing interests: The authors have no conflict of interest or financial interests for this article.
  • Ethics approval: The submitted work is original and has not been submitted to more than one journal for simultaneous consideration.
  • Availability of data and materials: This manuscript has no associated data.
  • Authors’ contributions: Conceptualization, methodology, investigation, validation, writing-original draft, review, editing, and reading have been performed by all authors of the paper.